### v1

# CryoGS: Gaussian Splatting for Cryo-EM Homogeneous Reconstruction

Abstract
As a critical modality for structural biology, cryogenic electron microscopy (cryo-EM) facilitates the determination of macromolecular structures at near-atomic resolution. The core computational task in single-particle cryo-EM is to reconstruct the 3D electrostatic potential of a molecule from a large collection of noisy 2D projections acquired at unknown orientations. Gaussian mixture models (GMMs) provide a continuous, compact, and physically interpretable representation for molecular density and have recently gained interest in cryo-EM reconstruction. However, existing methods rely on external consensus maps or atomic models for initialization, limiting their use in self-contained pipelines. Addressing this issue, we introduce cryoGS, a GMM-based method that integrates Gaussian splatting with the physics of cryo-EM image formation. In particular, we develop an orthogonal projection-aware Gaussian splatting, with adaptations such as a normalization term and FFT-aligned coordinate system tailored for cryo-EM imaging. All these innovations enable stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images using random initialization. Experimental results on real datasets validate the effectiveness and robustness of cryoGS over representative baselines. The code will be released upon publication.
footnotetext: Preprint. Under review.
Introduction
Single particle cryogenic electron microscopy (cryo-EM) has emerged as a transformative tool in structural biology, enabling visualization of macromolecular complexes at atomic or near-atomic resolution without crystallization (Kühlbrandt 2014; Nogales 2016; Renaud et al. 2018). Central to cryo-EM is the computational task of reconstructing a 3D volume from a large collection of 2D projection images, each corresponding to a different, unknown viewing direction of identical particles embedded in vitreous ice.
This inverse problem is inherently ill-posed and poses significant computational challenges. First, cryo-EM images are severely corrupted by noise due to the low electron dose required to prevent radiation damage. For experimental datasets, the signal-to-noise (SNR) could be as low as around 20-20 dB (Bendory, Bartesaghi, and Singer 2020; Bepler et al. 2020). Second, the orientations (poses) of individual particles are unknown and must be inferred jointly with the 3D structure. Third, many biological samples exhibit structural heterogeneity, with multiple conformational states coexisting in the same dataset.
These difficulties underscore two central objectives in cryo-EM image analysis: ab initio reconstruction, which aims to estimate both the 3D structure and particle poses directly from raw data without prior models, and heterogeneous reconstruction, which seeks to disentangle and reconstruct multiple structural states from the dataset. Both objectives fundamentally rely on the availability of a robust and efficient homogeneous reconstruction method, which assumes all particles correspond to a single structure and serves as a building block for more complex inference.
Approaches to homogeneous reconstruction include methods based on backprojection, iterative expectation-maximization with voxel-based volumes (Tang et al. 2007; Scheres 2012; Punjani et al. 2017; Shekarforoush et al. 2024), and more recently, neural representation learning (Zhong et al. 2021a, b), which models the 3D volume using coordinate-based networks. In parallel, Gaussian mixture models (GMMs) have received attention for their continuous, compact, and physically interpretable parameterization of molecular density (Chen and Ludtke 2021; Chen, Schmid, and Chiu 2023). Notably, GMMs offer a natural connection to atomic models and can represent fine structural details using fewer parameters (Chen, Toader, and Roy 2023; Li et al. 2024; Schwab et al. 2024; Chen 2025).
Despite their conceptual appeal, existing GMM-based methods for cryo-EM reconstruction rely on nontrivial prerequisite steps. These approaches (2021; 2023; 2023; 2024; 2024; 2025) typically rely on consensus volumes from external pipelines, or even atomic models, for initialization, and have not demonstrated stable convergence when directly optimizing from experimental images. In fact, no prior method achieves reliable GMM-based reconstruction even under known particle poses, due to the inherent difficulty of optimizing mixture parameters in extreme noise. As a result, GMMs lack a self-contained and stable formulation that can serve as a backbone for broader reconstruction workflows.
In this work, we propose cryoGS, a GMM-based homogeneous reconstruction method that fills this foundational gap. Given known particle poses, cryoGS performs stable and efficient reconstruction directly from raw cryo-EM projection images, starting from random initialization and requiring no external priors. Inspired by recent advances in 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), we model the 3D density as a mixture of anisotropic Gaussians and project them into 2D using a novel differentiable orthographic splatting algorithm consistent with cryo-EM physics. To support practical scalability and training efficiency, we develop a CUDA-accelerated real-space renderer that enables fast rasterization and optimization of the GMM.
Our contributions can be summarized as follows:
  • A self-contained GMM-based reconstruction method: We present cryoGS as the first method capable of performing cryo-EM homogeneous reconstruction from scratch using only a Gaussian Mixture Model, thereby establishing the missing foundation needed to develop GMMs into standalone reconstruction tools.
  • A physically accurate projection model: We design a splatting algorithm under orthogonal projection tailored to cryo-EM image formation, enabling differentiable projection of anisotropic Gaussians in real space.
  • An efficient implementation: We implement a CUDA-accelerated real-space renderer that enables fast optimization of GMMs with tens of thousands of Gaussians.
  • Experimental validation: We demonstrate the effectiveness of cryoGS on real datasets, showing it converges reliably from random initialization and achieves reconstruction quality outperforming state-of-the-art methods.
Related Work
Volume Representation in Cryo-EM
In cryo-EM experiments, purified biomolecules are rapidly frozen in a thin layer of vitreous ice, where each particle adopts a random orientation. A high-energy electron beam passes through the specimen, interacts with the electrostatic potential of the particles, and is recorded on a detector as a 2D projection image (Singer and Sigworth 2020). The goal of cryo-EM reconstruction is to recover the 3D electrostatic potential, i.e., the volume, from a large set of such noisy and randomly oriented 2D projections. Central to cryo-EM reconstruction is the choice of volume representation.
Voxel-based representations
are the most widely used in conventional cryo-EM software, e.g., RELION (2016), cryoSPARC (2017) and EMAN2 (2007). The 3D volume is discretized into a regular grid of density values, enabling fast projection and reconstruction via FFT-based algorithms. Despite their practical success, voxel grids are memory-intensive and inherently discrete, which limits their compatibility with modern learning-based analysis frameworks.
Neural fields
represent the volume as a continuous function parameterized by neural networks. These methods (Zhong et al. 2021a, b; Levy et al. 2022a, b, 2025) offer differentiability, implicit smoothness, and natural compatibility with learning-based heterogeneous analysis. However, the implicit nature of neural fields often comes at the cost of interpretability, and such models are typically slow to train and difficult to constrain with biological priors.
Gaussian mixture models
have a long history in structural biology, with early uses for molecular approximation (Grant and Pickup 1995; Grant, Gallardo, and Pickup 1996; Kawabata 2008). E2GMM (Chen and Ludtke 2021) was among the first to apply GMMs to cryo-EM heterogeneous reconstruction. Like neural fields, GMMs can approximate any smooth density function and support differentiable optimization. More importantly, GMMs provide an explicit and interpretable representation that naturally links to atomic structures. Recent studies (Chen, Schmid, and Chiu 2023; Chen, Toader, and Roy 2023; Li et al. 2024; Schwab et al. 2024; Chen 2025) have shown that GMMs can capture molecular flexibility by modeling atomic motion directly, making them highly suitable for heterogeneous reconstruction and downstream structural analysis.
However, existing GMM-based methods typically require initialization from an externally reconstructed consensus map or even an atomic model. Without such guidance, random initialization leads to unstable optimization and poor reconstruction quality. Our work addresses this limitation by introducing a GMM-based reconstruction pipeline that can be stably trained from scratch.
Gaussian Splatting
3DGS (Kerbl et al. 2023) is a recent differentiable rendering technique developed for real-time novel view synthesis. It represents a 3D scene as a collection of anisotropic Gaussians and renders images via rasterization-based accumulation and alpha blending (Zwicker et al. 2002). While 3DGS achieves high visual fidelity in synthetic and real-world RGB datasets, as a volume rendering method, it is not a physically accurate model of natural image formation (Huang et al. 2024).
Although the original 3DGS formulation is not physically consistent with natural image formation, its volume rendering framework closely aligns with the cryo-EM imaging model, where each image arises from an orthographic line integral of electrostatic potential modulated by the contrast transfer function (CTF). Leveraging this alignment, we adapt splatting to cryo-EM by rederiving the projection of anisotropic Gaussians under cryo-EM physics, replacing heuristic alpha blending with physically accurate line integrals and incorporating CTF modulation.
Method
Figure 1: Cryo-EM reconstruction aims to recover a 3D volume (a) from a large set of 2D particle images (b). (b) Purified biomolecules with random orientations are embedded in a thin layer of vitreous ice. The electrostatic potential of the sample interacts with transmitted electrons, forming a micrograph that contains 2D projections of the molecules. Individual particle images are extracted from the micrograph; they are extremely noisy and modulated by highly oscillatory CTFs. (c) Fourier slice theorem: the 2D Fourier transform of a particle image corresponds to a central slice of the 3D Fourier transform of the volume. (d) Overview of cryoGS. An anisotropic GMM representing the 3D volume is transformed to the projection direction, orthogonally projected onto a 2D image plane using a fast differentiable rasterizer, and modulated by the oscillatory CTF to simulate a physically accurate projection. The GMM parameters are optimized via gradients propagated from the discrepancy between the simulated and observed particle images. The resulting GMM can be voxelized to obtain the final 3D volume.
Overview
Our goal is to achieve physically accurate and computationally efficient cryo-EM reconstruction by leveraging a Gaussian Mixture Model (GMM). To this end, we propose cryoGS, a differentiable framework that represents the 3D electrostatic potential of a specimen as a set of anisotropic Gaussians and directly simulates the cryo-EM image formation process in real space, faithfully adhering to the physics of transmission electron microscopy.
Building upon recent progress in differentiable volume rendering, particularly the Gaussian splatting framework introduced in 3DGS (2023), we adopt a tile-based rasterization strategy to enable scalable and efficient computation. However, the original 3DGS formulation is not directly applicable to cryo-EM due to several fundamental mismatches: (i) it employs perspective projection consistent with a pinhole camera model, in contrast to the orthographic projection in cryo-EM imaging; (ii) it prioritizes photorealistic novel view synthesis over physical reconstruction accuracy; and (iii) its coordinate system is misaligned with the Fourier slice theorem, which underpins cryo-EM reconstruction.
To address these issues, cryoGS introduces several key adaptations: (i) we replace the heuristic alpha compositing with physically grounded line integrals to reflect the transmission nature of electron imaging; (ii) we fix the normalization between 3D-to-2D transformation and apply consistent learning rates across all parameters to ensure stable optimization; and (iii) we align the rasterization coordinate system with the FFT grid, allowing accurate gradient propagation through CTF modulation.
These modifications collectively enable cryoGS to perform stable, end-to-end differentiable reconstruction from raw cryo-EM particle images, starting from random initialization without relying on externally provided volumes.
Image Formation
As shown in Fig. 1(b), a beam of electrons passes through a vitrified specimen, and the transmitted wavefront undergoes phase shifts due to the specimen’s electrostatic potential(Singer and Sigworth 2020). Under the weak phase approximation, these phase shifts are linearly related to the 3D potential (volume), and the image formed at the detector is a line integral (projection) of this potential along the beam direction, further convolved with $\mathcal{H}$, the point spread function (PSF) of the imaging system.
In homogeneous reconstruction, it is assumed that all particle images $\mathcal{I}$ correspond to identical copies of a single 3D volume $\mathcal{V}:\mathbb{R}^{3}\rightarrow\mathbb{R}$, and that any conformational or compositional heterogeneity is negligible. Under this assumption, the image formation model can be expressed as:
$\mathcal{I}(r_{x},r_{y})=\mathcal{H}(r_{x},r_{y})*\int_{\mathbb{R}}\mathcal{V}(\mathbf{W}^{\top}\mathbf{r}+\mathbf{t})dr_{z}+\epsilon,$ (1)
where $\mathbf{r}=[r_{x},r_{y},r_{z}]^{\top}$ denotes the 3D Cartesian coordinates in real space, $\mathbf{W}\in SO(3)$ is the 3D pose of the volume, and $\mathbf{t}=[t_{x},t_{y},0]^{\top}$ is the in-plane translation, accounting for imperfect centering during particle cropping. The noise term $\epsilon$ is typically modeled as independent, zero-centered Gaussian noise.
According to the Fourier slice theorem (Bracewell 1956), illustrated in Fig. 1(c), the 2D Fourier transform of a projection corresponds to a central slice of the 3D Fourier transform of the volume, orthogonal to the projection direction and passing through the origin. Based on this property, an alternative and widely adopted formulation models reconstruction directly in the Fourier domain, where the image formation model becomes:
$\widehat{\mathcal{I}}(k_{x},k_{y})=\widehat{\mathcal{H}}(k_{x},k_{y})\cdot\widehat{\mathcal{V}}(\mathbf{W}^{\top}\mathbf{k})\cdot e^{-2\pi i\mathbf{k}^{\top}\mathbf{t}}+\widehat{\epsilon}$, (2)
where $\mathbf{k}=[k_{x},k_{y},0]^{\top}$ denotes the Cartesian coordinates in Fourier space, and the 2D spectrum $\widehat{\mathcal{I}}$, the CTF $\widehat{\mathcal{H}}$ and the 3D spectrum $\widehat{\mathcal{V}}$ denote the Fourier transform of $\mathcal{I}$, $\mathcal{H}$ and $\mathcal{V}$, respectively. The noise term $\widehat{\epsilon}$ is similarly modeled as independent, zero-mean Gaussian noise in the Fourier domain.
In this work, departing from most existing approaches that adopt Eq. (2), we instead build our pipeline on Eq. (1), performing homogeneous reconstruction directly in real space.
CryoGS
Anisotropic GMMs
are developed to represent the volume, which can be written in the form
$\mathcal{V}(\mathbf{r})=\sum_{i=1}^{N}A_{i}\mathcal{G}_{i}(\mathbf{r}),$ (3)
where $N$ denotes the number of Gaussians and $A_{i}$ is the amplitude of the $i$-th normalized Gaussian $\mathcal{G}_{i}(\mathbf{r})$.
By substituting Eq. (1), we obtain the full forward process of cryoGS. Specifically, we apply a viewing transformation to align the GMM to the target orientation, orthographically project each Gaussian along the z-axis to form a 2D image, and convolve the result with the PSF:
$\mathcal{I}(r_{x},r_{y})=\mathcal{H}(r_{x},r_{y})*\sum_{i=1}^{N}A_{i}\int_{\mathbb{R}}\mathcal{G}_{i}(\mathbf{W}^{\top}\mathbf{r}+\mathbf{t})dr_{z}.$ (4)
Since the integral is linear, each Gaussian contributes independently to the final image. We thus focus on a single Gaussian and omit the subscript $i$ in the following discussion. A normalized 3D Gaussian is defined as:
$\mathcal{G}(\mathbf{r}|\boldsymbol{\mu},\boldsymbol{\Sigma})=\frac{1}{(2\pi)^{\frac{3}{2}}|\boldsymbol{\Sigma}|^{\frac{1}{2}}}\exp(-\frac{1}{2}(\mathbf{r}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}(\mathbf{r}-\boldsymbol{\mu})),$ (5)
where $\boldsymbol{\mu}\in\mathbb{R}^{3}$ and $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$ denote the mean (center) and the covariance matrix, respectively. The determinant $|\boldsymbol{\Sigma}|$ ensures proper normalization. Following (Kerbl et al. 2023), to guarantee the positive semidefinite property, we construct the covariance matrix as:
$\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top},$ (6)
where $\mathbf{S}=\mathrm{diag}(\mathbf{s})$ is a diagonal scaling matrix and $\mathbf{R}\in SO(3)$ is a rotation matrix. In our implementation, we store the scaling vector $\mathbf{s}=[s_{x},s_{y},s_{z}]^{\top}$ and parameterize $\mathbf{R}$ using a quaternion $\mathbf{q}=[q_{w},q_{x},q_{y},q_{z}]^{\top}$. To ensure positivity and stable gradients during optimization, we apply a softplus function to both the amplitude $A$ and the scaling vector $\mathbf{s}$. The quaternion $\mathbf{q}$ is normalized to ensure it represents a valid rotation. Altogether, each anisotropic Gaussian is parameterized by the 11-dimensional set $\{\mu_{x},\mu_{y},\mu_{z},s_{x},s_{y},s_{z},q_{w},q_{x},q_{y},q_{z},A\}$.
When the scaling is isotropic, i.e., $s_{x}=s_{y}=s_{z}=\sigma$, the anisotropic Gaussian reduces to the standard isotropic form:
$\mathcal{G}(\mathbf{r}|\boldsymbol{\mu},\sigma)=\frac{1}{(2\pi)^{\frac{3}{2}}\sigma}\exp(-\frac{\|\mathbf{r}-\boldsymbol{\mu}\|_{2}^{2}}{2\sigma^{2}}),$ (7)
which is more commonly used in previous cryo-EM reconstruction works (2021; 2023; 2023; 2024; 2024; 2025).
The viewing transformation
is the first step in simulating image formation, aligning each Gaussian with a given projection direction. Since the parameters $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$ describe Gaussians in world coordinates, we must transform them into the image-space coordinates before projection.
According to the derivation in (Zwicker et al. 2002), applying an affine transformation to a Gaussian results in another Gaussian with appropriately transformed parameters. In our case, the transformation consists of a rotation $\mathbf{W}\in SO(3)$ and a 2D in-plane translation $\mathbf{t}\in\mathbb{R}^{3}$, leading to:
$\dot{\mathcal{G}}(\mathbf{r}|\dot{\boldsymbol{\mu}},\dot{\boldsymbol{\Sigma}})=\mathcal{G}(\mathbf{W}^{\top}\mathbf{r}+\mathbf{t}|\boldsymbol{\mu},\boldsymbol{\Sigma}),$ (8)
where the transformed mean and covariance are given by $\dot{\boldsymbol{\mu}}=\mathbf{W}\boldsymbol{\mu}+\mathbf{t}$ and $\dot{\boldsymbol{\Sigma}}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}$, respectively.
The orthogonal projection
[…]
$\tilde{\mathcal{G}}(\tilde{\mathbf{r}}|\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}})=\int_{\mathbb{R}}\dot{\mathcal{G}}(\mathbf{r}|\dot{\boldsymbol{\mu}},\dot{\boldsymbol{\Sigma}})dr_{z}.$ (9)
This operation effectively integrates the 3D Gaussian along the projection axis, preserving its Gaussian form in 2D. The resulting closed-form expression is:
$),\tilde{\mathcal{G}}(\tilde{\mathbf{r}}|\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}})=\frac{1}{2\pi|\tilde{\boldsymbol{\Sigma}}|^{\frac{1}{2}}}\exp(-\frac{1}{2}(\tilde{\mathbf{r}}-\tilde{\boldsymbol{\mu}})^{\top}\tilde{\boldsymbol{\Sigma}}^{-1}(\tilde{\mathbf{r}}-\tilde{\boldsymbol{\mu}}$)), (10)
where $\tilde{\mathbf{r}}=[r_{x},r_{y}]^{\top}$ denotes the 2D Cartesian coordinates in real space. As discussed in (Zwicker et al. 2002), this projection can be computed analytically by removing the z-axis components from the mean and covariance:
$\begin{cases}\dot{\boldsymbol{\mu}}=[\dot{\mu}_{x},\dot{\mu}_{y},\dot{\mu}_{z}]^{\top}\Rightarrow[\dot{\mu}_{x},\dot{\mu}_{y}]^{\top}=\tilde{\boldsymbol{\mu}}\\ \dot{\boldsymbol{\Sigma}}=\begin{bmatrix}\dot{\sigma}_{xx}&\dot{\sigma}_{xy}&\dot{\sigma}_{xz}\\ \dot{\sigma}_{xy}&\dot{\sigma}_{yy}&\dot{\sigma}_{yz}\\ \dot{\sigma}_{xz}&\dot{\sigma}_{yz}&\dot{\sigma}_{zz}\\ \end{bmatrix}\Rightarrow\begin{bmatrix}\dot{\sigma}_{xx}&\dot{\sigma}_{xy}\\ \dot{\sigma}_{xy}&\dot{\sigma}_{yy}\\ \end{bmatrix}=\tilde{\boldsymbol{\Sigma}}\end{cases}$ (11)
In prior works, such as 3DGS, the normalization term $1/(2\pi|\tilde{\boldsymbol{\Sigma}}|^{\frac{1}{2}})$ is often omitted, as their primary focus is on photorealistic novel view synthesis rather than the physical fidelity of the underlying 3D representation. However, in cryo-EM reconstruction, the ultimate goal is to recover the correct 3D volume. Omitting this normalization introduces bias in amplitude and leads to incorrect reconstructions. Therefore, unlike 3DGS, we retain the normalization term to preserve the quantitative correctness of the model.
After projection, the final image is constructed by summing the weighted contributions of all projected 2D Gaussians and applying the PSF:
[…]
In practice, consistent with typical cryo-EM pipelines, we apply the CTF instead in the Fourier domain:
[…]
where $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the Fourier and inverse Fourier transform operators, respectively.
Fast differentiable rasterization.
We adopt the efficient tile-based rasterization framework from 3DGS (Kerbl et al. 2023), which enables scalable and differentiable processing of tens of thousands of Gaussians via per-tile accumulation. Unlike 3DGS, which uses alpha blending for photorealistic rendering, we modify the rasterization to directly sum projected Gaussian contributions, in accordance with the physical transmission model in cryo-EM.
A key reason we choose to operate in real space is that it allows us to fully exploit this fast rasterization strategy. In high-resolution reconstructions, individual Gaussians in real space have small spatial scales and affect only a few nearby pixels. This locality means each GPU thread—responsible for a single pixel—only needs to process a small subset of all Gaussians. In contrast, Gaussians in Fourier space become broad as resolution increases, leading to near-global support. As a result, each pixel in the frequency domain must aggregate contributions from nearly all Gaussians, making fast rendering impractical.
For an image of shape $D\times D$, the original 3DGS implementation places the continuous coordinate center at $(D-1)/2,(D-1)/2]^{\top}$, i.e., halfway between two discrete pixels. In contrast, FFT-based image formation assumes the origin is located at the integer grid point $\lfloor D/2\rfloor,\lfloor D/2\rfloor]^{\top}$. To ensure compatibility with FFT-based forward and backward modeling, we shift the rasterization coordinates by half a pixel so that the image center aligns with the FFT grid.
This alignment eliminates phase inconsistencies and enables accurate electron projection simulation, while preserving the computational efficiency of the 3DGS architecture.
Loss function.
Unlike previous GMM-based methods, which often rely on specially designed loss functions combined with complex regularization or constraints to ensure stable optimization, we adopt a much simpler formulation. Specifically, we directly apply the mean squared error (MSE) loss between the GMM-based projection $\mathcal{I}$ and the observed image $\mathcal{I}_{\rm ref}$:
$\mathcal{L}=\|\mathcal{I}-\mathcal{I}_{\rm ref}\|_{2}.$ (14)
Despite its simplicity, this loss formulation leads to stable and fast convergence in practice, without requiring additional regularization terms.
Experiment
Figure 2: Qualitative and quantitative comparison of voxel-based, neural, and GMM-based representations. (Left) Final 3D reconstructions on four real datasets visualized with ChimeraX (Pettersen et al. 2021). (Right) FSC curves are plotted for quantitative evaluation. Gray dashed lines indicate the standard resolution thresholds of $0.5$ and $0.143$, reported in Angstroms (Å). CryoGS consistently achieves higher resolution across all datasets.
Figure 3: Evaluation of reconstruction performance with varying numbers of Gaussians. Increasing the number improves both accuracy and robustness, with 10,00010{,}000 Gaussians generally sufficient for most datasets.
Figure 4: Runtime efficiency comparison across reconstruction methods at different resolutions ($D=192$ and $D=256$). Frame rates (FPS) are measured under increasing numbers of Gaussians (log-scaled).
Experimental Settings
Datasets.
We evaluate our method on four publicly available cryo-EM datasets from the Electron Microscopy Public Image Archive (EMPIAR) (Iudin et al. 2016): EMPIAR-10028 (Pf80S ribosome) (Wong et al. 2014), EMPIAR-10049 (RAG complex) (Ru et al. 2015), EMPIAR-10076 (E. coli LSU assembly) (Davis et al. 2016), and EMPIAR-10180 (pre-catalytic spliceosome) (Plaschka, Lin, and Nagai 2017). These datasets span a range of structural complexity and image quality, from rigid assemblies with high contrast to highly heterogeneous macromolecular machines. For each dataset, we use the provided particle images, consensus pose estimates, and CTF parameters. All reconstructions are performed under the homogeneous assumption.
Evaluation metrics.
Since ground truth volumes are unavailable for real datasets, we follow standard practice and assess reconstruction quality using the gold standard Fourier Shell Correlation (FSC) (Van Heel and Schatz 2005). Each dataset is split evenly into two halves, and the method is applied independently to each subset. The resulting reconstructions are compared in Fourier space by computing FSC as a function of spatial frequency, which quantifies their consistency across frequency shells. The resolution is defined as the spatial frequency at which the FSC curve drops below the $0.143$ threshold.
Implementation details.
For all experiments, particle images from EMPIAR-10028, 10076, and 10180 are downsampled to $256\times 256$, while EMPIAR-10049 is used at its original $192\times 192$ resolution. Published particle translations are applied to the observed images via phase shifting in Fourier space prior to reconstruction, rather than through the GMM viewing transform. The 3D volume domain is defined as $0.5,0.5]^{3}$, assuming each projection lies within $0.5,0.5]^{2}$ in the image plane. Gaussians are initialized with random means $\boldsymbol{\mu}\sim\mathcal{N}(\mathbf{0},0.075^{2}\mathbf{I})$, isotropic scales $s_{x}\!=\!s_{y}\!=\!s_{z}\!=\!0.0075$, identity quaternion $\mathbf{q}\!=\![1,0,0,0]^{T}$, and amplitude $A=1/(2N)$, where $N$ is the number of Gaussians. All parameters are trainable. We use Adam (Kingma and Ba 2014) with batch size $1$, learning rate $0.001$, and exponential decay ($\gamma=0.1$) at each epoch. All GMMs are trained for $5$ epochs. For the voxel-based backprojection baseline, we use cryoDRGN’s official implementation (Zhong et al. 2021a). For neural representation learning, we follow cryoDRGN’s default configuration: three 1,0241{,}024-node layers, trained for $50$ epochs. Since related methods (2021a; 2021b; 2022a; 2022b; 2025) share similar neural fields and differ mainly in pose estimation, we focus comparison on cryoDRGN. All experiments are run on a single NVIDIA GeForce RTX 3090.
Evaluation on Real Datasets
We evaluate the performance of different volume representations on real cryo-EM datasets under a homogeneous reconstruction setting. To ensure a fair comparison focused solely on the choice of volume representation, all methods reconstruct consensus maps using the same set of published particle poses, without performing pose estimation. Our evaluation focuses on two aspects: (i) the ability to reconstruct high-resolution consensus maps, and (ii) robustness to noise and imperfect pose assignments. We compare voxel-based backprojection, the neural representation method cryoDRGN (Zhong et al. 2021a), and our proposed GMM-based method cryoGS. Visualizations of the reconstructed volumes are shown in Fig. 2, and spatial resolution is quantified using gold-standard FSC curves. Each volume of cryoGS is represented using 30,00030{,}000 Gaussians.
EMPIAR-10028.
The Pf80S ribosome is relatively easy to reconstruct due to its high-contrast images and structurally stable particles. All methods achieve high-resolution results ($3.80$ Å) and strong FSC agreement across the spectrum. cryoDRGN yields slightly higher FSC values at intermediate frequencies, while cryoGS outperforms all baselines at high spatial frequencies, demonstrating its ability to recover fine structural details.
EMPIAR-10049.
The RAG complex poses greater challenges due to imperfect published poses and flexible regions such as the DNA elements and the nonamer binding domain (NBD), indicated by arrows. CryoGS significantly outperforms both backprojection and cryoDRGN, achieving a resolution of $2.49$ Å. Unlike the baselines, cryoGS reconstructs the DNA elements and the NBD with minimal artifacts. Its FSC curve remains consistently above those of other methods across all spatial frequencies, highlighting its robustness to pose inaccuracies and structural variability.
EMPIAR-10076.
This LSU assembly dataset contains substantial compositional and conformational heterogeneity, making consensus reconstruction particularly challenging. Both FSC analysis and visualization show that cryoGS is more resilient under such conditions, achieving a resolution of $3.30$ Å with fewer artifacts than voxel-based or neural methods, as indicated by the red circle.
EMPIAR-10180.
The spliceosome dataset features large-scale motions of the SF3b indicated by the red circle, making consensus reconstruction particularly challenging. The reconstructions from backprojection and cryoDRGN show pronounced artifacts in this region, while cryoGS is more robust to such motions and achieves a resolution of $4.26$ Å. FSC analysis further confirms that cryoGS significantly outperforms the baselines across the frequency range.
CryoGS consistently converges within just $5$ epochs, with FSC curves from the $4$th and $5$th epochs tightly overlapping, indicating stable optimization and improved generalization in consensus reconstruction.
Ablation Studies
Number of Gaussians.
Fig. 3 shows the FSC curves for cryoGS with varying numbers of Gaussians. In general, increasing the number of Gaussians leads to improved FSC, as a denser GMM provides greater representational capacity. While cryoGS performs well on most datasets, its relative performance varies due to differences in structural complexity and dataset-specific challenges. On EMPIAR-10028, cryoGS reaches a resolution of $3.8$ Å under all settings. While configurations with fewer than 10,00010{,}000 Gaussians exhibit lower FSC values than cryoDRGN and backprojection across most frequencies, the curves intersect at the highest frequency, indicating comparable final resolution. For EMPIAR-10049, all cryoGS settings significantly outperform both backprojection ($4.00$ Å) and cryoDRGN ($4.07$ Å), achieving a resolution of $2.49$ Å. Moreover, the FSC curves of all cryoGS variants remain consistently above those of the two baselines across the entire frequency range. For EMPIAR-10076, the 30,00030{,}000-Gaussian model clearly outperforms other settings; even with fewer Gaussians, cryoGS still surpasses the baselines, reaching $3.3$ Å. For EMPIAR-10180, the models with 10,00010{,}000 and 30,00030{,}000 Gaussians achieve the best FSC, reaching $4.3$ Å, while sparser GMMs remain competitive at high spatial frequencies. Overall, we find that using 10,00010{,}000 Gaussians is sufficient to reconstruct high-resolution volumes that consistently outperform baseline methods across most datasets.
Runtime efficiency.
[…]
Conclusion
We present cryoGS, a novel GMM-based framework that integrates Gaussian splatting with the physics of cryo-EM image formation. CryoGS enables stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images, starting from random initialization without relying on consensus volumes. Experimental results on real datasets demonstrate the effectiveness, robustness, and faster convergence of cryoGS compared to representative baselines.
Limitation and future work. While our current method assumes known poses and thus does not qualify as an ab initio approach, cryoGS establishes a principled foundation for future GMM-based methods that aim to tackle ab initio and heterogeneous reconstruction. We believe cryoGS provides a missing piece in the broader goal of integrating GMMs into scalable and self-contained cryo-EM reconstruction pipelines. These directions are left for future work.
Appendix
Appendix A Dataset Details
We provide detailed statistics and characteristics of the four cryo-EM datasets used in our experiments:
  • EMPIAR-10028 (Plasmodium falciparum 80S (Pf80S) ribosome) (Wong et al. 2014): 105,247105{,}247 particle images of size $360\times 360$ pixels at a sampling rate of $1.34$ Å/pixel. This is a widely used benchmark with high-contrast images and a static structure.
  • EMPIAR-10049 (RAG1-RAG2 complex) (Ru et al. 2015): 108,544108{,}544 particles of size $192\times 192$ pixels at $1.23$ Å/pixel. This dataset is considered more challenging due to its lower contrast and flexibility in some regions.
  • EMPIAR-10076 (E. coli large ribosomal subunit undergoing (LSU) assembly) (Davis et al. 2016): 131,899131{,}899 particles of size $320\times 320$ pixels at $1.31$ Å/pixel. This dataset contains substantial conformational and compositional heterogeneity, which poses a challenge to homogeneous modeling.
  • EMPIAR-10180 (Pre-catalytic spliceosome) (Plaschka, Lin, and Nagai 2017): 327,490327{,}490 particles of size $320\times 320$ pixels at $1.69$ Å/pixel. It samples a continuum of conformations, particularly involving large-scale motions of the SF3b subcomplex.
Appendix B Fourier Shell Correlation
To evaluate reconstruction quality on real datasets where ground truth volumes are unavailable, we adopt the gold standard Fourier Shell Correlation (FSC) (Van Heel and Schatz 2005), following established protocols. Each dataset is randomly split into two halves, and the reconstruction algorithm is applied independently to each subset. Let the resulting volumes be $\widehat{\mathcal{V}}_{a}(\mathbf{k})$ and $\widehat{\mathcal{V}}_{b}(\mathbf{k})$, representing their Fourier transforms. The FSC is computed as a function of spatial frequency $k=|\mathbf{k}|$ using the following formula:
$\mathrm{FSC}(k)=\frac{\sum\limits_{|\mathbf{k}|=k}\widehat{\mathcal{V}}_{a}(\mathbf{k})\cdot\widehat{\mathcal{V}}_{b}(\mathbf{k})^{*}}{\sqrt{\Big{(}\sum\limits_{|\mathbf{k}|=k}|\widehat{\mathcal{V}}_{a}(\mathbf{k})|^{2}\Big{)}\Big{(}\sum\limits_{|\mathbf{k}|=k}|\widehat{\mathcal{V}}_{b}(\mathbf{k})|^{2}\Big{)}}}.$ (B.1)
This metric quantifies the correlation between two independently reconstructed volumes within concentric shells in Fourier space. The spatial resolution is defined as the frequency where the FSC curve drops below the $0.143$ threshold, indicating the limit of reproducible structural detail.
Appendix C More Implementation Details
For all experiments, we do not apply any windowing to the observed particle images during preprocessing. The 3D volume is defined over the domain $E,E]^{3}$, and each 2D projection is assumed to lie within $E,E]^{2}$ in the image plane, where $E=0.5$ defines the spatial extent.
Intuition Behind Initialization
These values are fixed but grounded in straightforward statistical intuition. We observe that most particles are concentrated within a spherical region of radius $E/2$. To reflect this prior and accelerate convergence, we initialize the Gaussian means within this region.
Moreover, based on the three-sigma rule for Gaussian distributions $\mathcal{N}(\mu,\sigma^{2})$, where $99.7\%$ of samples fall within $\mu-3\sigma,\mu+3\sigma]$, we obtain $\sigma=E/6$ from $3\sigma=E/2$. To slightly tighten the spread, we apply a scaling factor and use $\sigma=0.9\cdot E/6=0.075$ to initialize the means. The initial scale of each Gaussian is set to $0.1\times 0.075=0.0075$, encouraging localized support. Finally, to maintain consistent overall energy across varying numbers of Gaussians, we initialize the amplitude as $A=1/(2N)$, where $N$ is the total number of Gaussians.
Methods # Params Settings Batch Size GPU Mem. (MiB)
$D=192$ $D=256$
Backprojection $(D+1)^{3}$ 11 $508$ $396$
CryoDRGN (2021a) $(6\cdot\lfloor D/2\rfloor+L+3)\cdot C$ C=1,024C=1{,}024 11 $680$ 1,0081{,}008
$L\cdot C^{2}+2$ $L=3$ 88 2,5602{,}560 4,9064{,}906
CryoGS (Ours) $11\cdot N$ N=2,048N=2{,}048 11 $344$ $346$
N=3,072N=3{,}072 $344$ $348$
N=5,120N=5{,}120 $346$ $348$
N=10, [^0] N=10{,}000 $348$ $350$
N=30, [^0] N=30{,}000 $376$ $378$
Table S1: Comparison of GPU memory usage across reconstruction methods at $D=192$ and $D=256$. CryoGS achieves the lowest and most stable memory footprint, while cryoDRGN incurs significantly higher usage due to its deep neural architecture. All measurements were performed on a single NVIDIA RTX 3090, reflecting the memory usage during the reconstruction of a single full density map. Memory for voxelization and FSC-related half-map reconstruction is not included.
Intuition Behind Learning Rate
In the original 3DGS (Kerbl et al. 2023), different learning rates are assigned to different types of Gaussian parameters (means, scales, rotations, opacities). While this works well in novel view synthesis, it introduces instability in cryo-EM reconstruction. Let the full parameter vector be $\boldsymbol{\theta}=[\mu_{x},\mu_{y},\mu_{z},s_{x},s_{y},s_{z},q_{w},q_{x},q_{y},q_{z},A]^{\top}$. In gradient descent optimization, the direction of parameter updates is determined by the gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}$. Unequal learning rates distort this direction by scaling different components unequally, which can lead to divergence. We observe that such practice causes Gaussians to spread uncontrollably in early iterations and finally diverge. To avoid this, we adopt a single unified learning rate across all parameter types, preserving the intended descent direction and ensuring stable convergence.
Appendix D Additional Results
Memory Usage
Tab. S1 presents a comparison of GPU memory usage across different reconstruction methods. CryoDRGN (Zhong et al. 2021a) exhibits the highest memory footprint, exceeding $2.5$ GiB at $D=192$ and approaching $5$ GiB at $D=256$, primarily due to its deep neural decoder and a larger batch size of 8. Interestingly, backprojection consumes more memory at $D=192$ than at $D=256$, which may be attributed to implementation-specific factors such as padding overhead or kernel-level optimizations that favor power-of-two dimensions. This anomaly appears method-specific and does not reflect a general trend. In contrast, cryoGS demonstrates consistently low memory usage across all configurations. Even with as many as 30,00030{,}000 Gaussians, cryoGS maintains a GPU memory footprint below $380$ MiB, with negligible variation across resolutions. This efficiency underscores the scalability and suitability of cryoGS for large-scale or memory-constrained cryo-EM reconstruction scenarios.
Figure S1: Qualitative and quantitative comparison of isotropic and anisotropic GMMs (N=30, [^0] N=30{,}000) on four real datasets. FSC curves show that anisotropic Gaussians consistently achieve higher correlations across spatial frequencies, indicating improved reconstruction accuracy and robustness. Volume visualizations further reveal that anisotropic GMMs better recover fine structural details and directionally extended shapes, whereas isotropic Gaussians tend to fragment such regions.
Isotropic vs. Anisotropic
We investigate how the choice between isotropic and anisotropic Gaussians affects reconstruction quality. While cryoGS adopts anisotropic Gaussians to represent 3D volumes, it remains compatible with the isotropic formulation commonly used in prior work (2021; 2023; 2023; 2024; 2025). As shown in Fig. S1, anisotropic GMMs achieve higher FSC scores across spatial frequencies and produce sharper, more detailed structures. Subjectively, isotropic Gaussians struggle to capture directionally extended structures and are more prone to noise, often resulting in oversmoothing or distortion. This geometric rigidity may impair their ability to fit complex structures during early optimization, contributing to the unstable convergence observed with random initialization in prior methods. These results highlight the improved representational capacity and reconstruction robustness enabled by anisotropic modeling.
Figure S2: Qualitative evaluation of reconstruction performance with different numbers of Gaussians. Increasing the number of Gaussians leads to visibly improved reconstructions, with finer structural details and enhanced sharpness. Red arrows mark representative regions that highlight the qualitative differences for clearer comparison across settings.
Number of Gaussians
We present visual comparisons of reconstruction results using different numbers of Gaussians. As shown in Fig. S2, increasing the number of components yields progressively sharper and more detailed structures. These qualitative observations align with the quantitative improvements in FSC curves reported in the main paper. Red arrows highlight representative regions where the differences in reconstruction quality are especially pronounced, facilitating direct visual comparison across settings.

### v2

@@
-CryoGS: Gaussian Splatting for Cryo-EM Homogeneous Reconstruction
+CryoSplat: Gaussian Splatting for Cryo-EM Homogeneous Reconstruction
 Abstract
@@
 However, existing methods rely on external consensus maps or atomic models for initialization, limiting their use in self-contained pipelines.
-Addressing this issue, we introduce cryoGS, a GMM-based method that integrates Gaussian splatting with the physics of cryo-EM image formation.
+Addressing this issue, we introduce cryoSplat, a GMM-based method that integrates Gaussian splatting with the physics of cryo-EM image formation.
 In particular, we develop an orthogonal projection-aware Gaussian splatting, with adaptations such as a normalization term and FFT-aligned coordinate system tailored for cryo-EM imaging.
 All these innovations enable stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images using random initialization.
-Experimental results on real datasets validate the effectiveness and robustness of cryoGS over representative baselines.
+Experimental results on real datasets validate the effectiveness and robustness of cryoSplat over representative baselines.
 The code will be released upon publication.
@@
 As a result, GMMs lack a self-contained and stable formulation that can serve as a backbone for broader reconstruction workflows.
-In this work, we propose cryoGS, a GMM-based homogeneous reconstruction method that fills this foundational gap.
-Given known particle poses, cryoGS performs stable and efficient reconstruction directly from raw cryo-EM projection images, starting from random initialization and requiring no external priors.
+In this work, we propose cryoSplat, a GMM-based homogeneous reconstruction method that fills this foundational gap.
+Given known particle poses, cryoSplat performs stable and efficient reconstruction directly from raw cryo-EM projection images, starting from random initialization and requiring no external priors.
 Inspired by recent advances in 3D Gaussian Splatting (3DGS) (Kerbl et al.
@@
 Our contributions can be summarized as follows:
-  • A self-contained GMM-based reconstruction method: We present cryoGS as the first method capable of performing cryo-EM homogeneous reconstruction from scratch using only a Gaussian Mixture Model, thereby establishing the missing foundation needed to develop GMMs into standalone reconstruction tools.
+  • A self-contained GMM-based reconstruction method: We present cryoSplat as the first method capable of performing cryo-EM homogeneous reconstruction from scratch using only a Gaussian Mixture Model, thereby establishing the missing foundation needed to develop GMMs into standalone reconstruction tools.
   • A physically accurate projection model: We design a splatting algorithm under orthogonal projection tailored to cryo-EM image formation, enabling differentiable projection of anisotropic Gaussians in real space.
   • An efficient implementation: We implement a CUDA-accelerated real-space renderer that enables fast optimization of GMMs with tens of thousands of Gaussians.
-  • Experimental validation: We demonstrate the effectiveness of cryoGS on real datasets, showing it converges reliably from random initialization and achieves reconstruction quality outperforming state-of-the-art methods.
+  • Experimental validation: We demonstrate the effectiveness of cryoSplat on real datasets, showing it converges reliably from random initialization and achieves reconstruction quality outperforming state-of-the-art methods.
 Related Work
@@
 (c) Fourier slice theorem: the 2D Fourier transform of a particle image corresponds to a central slice of the 3D Fourier transform of the volume.
-(d) Overview of cryoGS.
+(d) Overview of cryoSplat.
 An anisotropic GMM representing the 3D volume is transformed to the projection direction, orthogonally projected onto a 2D image plane using a fast differentiable rasterizer, and modulated by the oscillatory CTF to simulate a physically accurate projection.
@@
 Our goal is to achieve physically accurate and computationally efficient cryo-EM reconstruction by leveraging a Gaussian Mixture Model (GMM).
-To this end, we propose cryoGS, a differentiable framework that represents the 3D electrostatic potential of a specimen as a set of anisotropic Gaussians and directly simulates the cryo-EM image formation process in real space, faithfully adhering to the physics of transmission electron microscopy.
+To this end, we propose cryoSplat, a differentiable framework that represents the 3D electrostatic potential of a specimen as a set of anisotropic Gaussians and directly simulates the cryo-EM image formation process in real space, faithfully adhering to the physics of transmission electron microscopy.
 Building upon recent progress in differentiable volume rendering, particularly the Gaussian splatting framework introduced in 3DGS (2023), we adopt a tile-based rasterization strategy to enable scalable and efficient computation.
 However, the original 3DGS formulation is not directly applicable to cryo-EM due to several fundamental mismatches: (i) it employs perspective projection consistent with a pinhole camera model, in contrast to the orthographic projection in cryo-EM imaging; (ii) it prioritizes photorealistic novel view synthesis over physical reconstruction accuracy; and (iii) its coordinate system is misaligned with the Fourier slice theorem, which underpins cryo-EM reconstruction.
-To address these issues, cryoGS introduces several key adaptations: (i) we replace the heuristic alpha compositing with physically grounded line integrals to reflect the transmission nature of electron imaging; (ii) we fix the normalization between 3D-to-2D transformation and apply consistent learning rates across all parameters to ensure stable optimization; and (iii) we align the rasterization coordinate system with the FFT grid, allowing accurate gradient propagation through CTF modulation.
-These modifications collectively enable cryoGS to perform stable, end-to-end differentiable reconstruction from raw cryo-EM particle images, starting from random initialization without relying on externally provided volumes.
+To address these issues, cryoSplat introduces several key adaptations: (i) we replace the heuristic alpha compositing with physically grounded line integrals to reflect the transmission nature of electron imaging; (ii) we fix the normalization between 3D-to-2D transformation and apply consistent learning rates across all parameters to ensure stable optimization; and (iii) we align the rasterization coordinate system with the FFT grid, allowing accurate gradient propagation through CTF modulation.
+These modifications collectively enable cryoSplat to perform stable, end-to-end differentiable reconstruction from raw cryo-EM particle images, starting from random initialization without relying on externally provided volumes.
 Image Formation
@@
 In this work, departing from most existing approaches that adopt Eq. (2), we instead build our pipeline on Eq. (1), performing homogeneous reconstruction directly in real space.
-CryoGS
+CryoSplat
 Anisotropic GMMs
@@
 By substituting Eq.
-(1), we obtain the full forward process of cryoGS.
+(1), we obtain the full forward process of cryoSplat.
 Specifically, we apply a viewing transformation to align the GMM to the target orientation, orthographically project each Gaussian along the z-axis to form a 2D image, and convolve the result with the PSF:
@@
 Gray dashed lines indicate the standard resolution thresholds of $0.5$ and $0.143$, reported in Angstroms (Å).
-CryoGS consistently achieves higher resolution across all datasets.
+CryoSplat consistently achieves higher resolution across all datasets.
 Figure 3: Evaluation of reconstruction performance with varying numbers of Gaussians.
@@
 We compare voxel-based backprojection, the neural representation method cryoDRGN (Zhong et al.
-2021a), and our proposed GMM-based method cryoGS.
+2021a), and our proposed GMM-based method cryoSplat.
 Visualizations of the reconstructed volumes are shown in Fig.
 2, and spatial resolution is quantified using gold-standard FSC curves.
-Each volume of cryoGS is represented using 30,00030{,}000 Gaussians.
+Each volume of cryoSplat is represented using 30,00030{,}000 Gaussians.
 EMPIAR-10028.
@@
 All methods achieve high-resolution results ($3.80$ Å) and strong FSC agreement across the spectrum.
-cryoDRGN yields slightly higher FSC values at intermediate frequencies, while cryoGS outperforms all baselines at high spatial frequencies, demonstrating its ability to recover fine structural details.
+cryoDRGN yields slightly higher FSC values at intermediate frequencies, while cryoSplat outperforms all baselines at high spatial frequencies, demonstrating its ability to recover fine structural details.
 EMPIAR-10049.
 The RAG complex poses greater challenges due to imperfect published poses and flexible regions such as the DNA elements and the nonamer binding domain (NBD), indicated by arrows.
-CryoGS significantly outperforms both backprojection and cryoDRGN, achieving a resolution of $2.49$ Å.
-Unlike the baselines, cryoGS reconstructs the DNA elements and the NBD with minimal artifacts.
+CryoSplat significantly outperforms both backprojection and cryoDRGN, achieving a resolution of $2.49$ Å.
+Unlike the baselines, cryoSplat reconstructs the DNA elements and the NBD with minimal artifacts.
 Its FSC curve remains consistently above those of other methods across all spatial frequencies, highlighting its robustness to pose inaccuracies and structural variability.
@@
 This LSU assembly dataset contains substantial compositional and conformational heterogeneity, making consensus reconstruction particularly challenging.
-Both FSC analysis and visualization show that cryoGS is more resilient under such conditions, achieving a resolution of $3.30$ Å with fewer artifacts than voxel-based or neural methods, as indicated by the red circle.
+Both FSC analysis and visualization show that cryoSplat is more resilient under such conditions, achieving a resolution of $3.30$ Å with fewer artifacts than voxel-based or neural methods, as indicated by the red circle.
 EMPIAR-10180.
 The spliceosome dataset features large-scale motions of the SF3b indicated by the red circle, making consensus reconstruction particularly challenging.
-The reconstructions from backprojection and cryoDRGN show pronounced artifacts in this region, while cryoGS is more robust to such motions and achieves a resolution of $4.26$ Å.
-FSC analysis further confirms that cryoGS significantly outperforms the baselines across the frequency range.
-CryoGS consistently converges within just $5$ epochs, with FSC curves from the $4$th and $5$th epochs tightly overlapping, indicating stable optimization and improved generalization in consensus reconstruction.
+The reconstructions from backprojection and cryoDRGN show pronounced artifacts in this region, while cryoSplat is more robust to such motions and achieves a resolution of $4.26$ Å.
+FSC analysis further confirms that cryoSplat significantly outperforms the baselines across the frequency range.
+CryoSplat consistently converges within just $5$ epochs, with FSC curves from the $4$th and $5$th epochs tightly overlapping, indicating stable optimization and improved generalization in consensus reconstruction.
 Ablation Studies
@@
 Fig.
-3 shows the FSC curves for cryoGS with varying numbers of Gaussians.
+3 shows the FSC curves for cryoSplat with varying numbers of Gaussians.
 In general, increasing the number of Gaussians leads to improved FSC, as a denser GMM provides greater representational capacity.
-While cryoGS performs well on most datasets, its relative performance varies due to differences in structural complexity and dataset-specific challenges.
-On EMPIAR-10028, cryoGS reaches a resolution of $3.8$ Å under all settings.
+While cryoSplat performs well on most datasets, its relative performance varies due to differences in structural complexity and dataset-specific challenges.
+On EMPIAR-10028, cryoSplat reaches a resolution of $3.8$ Å under all settings.
 While configurations with fewer than 10,00010{,}000 Gaussians exhibit lower FSC values than cryoDRGN and backprojection across most frequencies, the curves intersect at the highest frequency, indicating comparable final resolution.
-For EMPIAR-10049, all cryoGS settings significantly outperform both backprojection ($4.00$ Å) and cryoDRGN ($4.07$ Å), achieving a resolution of $2.49$ Å.
-Moreover, the FSC curves of all cryoGS variants remain consistently above those of the two baselines across the entire frequency range.
-For EMPIAR-10076, the 30,00030{,}000-Gaussian model clearly outperforms other settings; even with fewer Gaussians, cryoGS still surpasses the baselines, reaching $3.3$ Å.
+For EMPIAR-10049, all cryoSplat settings significantly outperform both backprojection ($4.00$ Å) and cryoDRGN ($4.07$ Å), achieving a resolution of $2.49$ Å.
+Moreover, the FSC curves of all cryoSplat variants remain consistently above those of the two baselines across the entire frequency range.
+For EMPIAR-10076, the 30,00030{,}000-Gaussian model clearly outperforms other settings; even with fewer Gaussians, cryoSplat still surpasses the baselines, reaching $3.3$ Å.
 For EMPIAR-10180, the models with 10,00010{,}000 and 30,00030{,}000 Gaussians achieve the best FSC, reaching $4.3$ Å, while sparser GMMs remain competitive at high spatial frequencies.
@@
 Conclusion
-We present cryoGS, a novel GMM-based framework that integrates Gaussian splatting with the physics of cryo-EM image formation.
-CryoGS enables stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images, starting from random initialization without relying on consensus volumes.
-Experimental results on real datasets demonstrate the effectiveness, robustness, and faster convergence of cryoGS compared to representative baselines.
+We present cryoSplat, a novel GMM-based framework that integrates Gaussian splatting with the physics of cryo-EM image formation.
+CryoSplat enables stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images, starting from random initialization without relying on consensus volumes.
+Experimental results on real datasets demonstrate the effectiveness, robustness, and faster convergence of cryoSplat compared to representative baselines.
 Limitation and future work.
-While our current method assumes known poses and thus does not qualify as an ab initio approach, cryoGS establishes a principled foundation for future GMM-based methods that aim to tackle ab initio and heterogeneous reconstruction.
-We believe cryoGS provides a missing piece in the broader goal of integrating GMMs into scalable and self-contained cryo-EM reconstruction pipelines.
+While our current method assumes known poses and thus does not qualify as an ab initio approach, cryoSplat establishes a principled foundation for future GMM-based methods that aim to tackle ab initio and heterogeneous reconstruction.
+We believe cryoSplat provides a missing piece in the broader goal of integrating GMMs into scalable and self-contained cryo-EM reconstruction pipelines.
 These directions are left for future work.
@@
 $L\cdot C^{2}+2$ $L=3$ 88 2,5602{,}560 4,9064{,}906
-CryoGS (Ours) $11\cdot N$ N=2,048N=2{,}048 11 $344$ $346$
+CryoSplat (Ours) $11\cdot N$ N=2,048N=2{,}048 11 $344$ $346$
 N=3,072N=3{,}072 $344$ $348$
@@
 Table S1: Comparison of GPU memory usage across reconstruction methods at $D=192$ and $D=256$.
-CryoGS achieves the lowest and most stable memory footprint, while cryoDRGN incurs significantly higher usage due to its deep neural architecture.
+CryoSplat achieves the lowest and most stable memory footprint, while cryoDRGN incurs significantly higher usage due to its deep neural architecture.
 All measurements were performed on a single NVIDIA RTX 3090, reflecting the memory usage during the reconstruction of a single full density map.
@@
 This anomaly appears method-specific and does not reflect a general trend.
-In contrast, cryoGS demonstrates consistently low memory usage across all configurations.
-Even with as many as 30,00030{,}000 Gaussians, cryoGS maintains a GPU memory footprint below $380$ MiB, with negligible variation across resolutions.
-This efficiency underscores the scalability and suitability of cryoGS for large-scale or memory-constrained cryo-EM reconstruction scenarios.
+In contrast, cryoSplat demonstrates consistently low memory usage across all configurations.
+Even with as many as 30,00030{,}000 Gaussians, cryoSplat maintains a GPU memory footprint below $380$ MiB, with negligible variation across resolutions.
+This efficiency underscores the scalability and suitability of cryoSplat for large-scale or memory-constrained cryo-EM reconstruction scenarios.
 Figure S1: Qualitative and quantitative comparison of isotropic and anisotropic GMMs (N=30, [^0] N=30{,}000) on four real datasets.
@@
 We investigate how the choice between isotropic and anisotropic Gaussians affects reconstruction quality.
-While cryoGS adopts anisotropic Gaussians to represent 3D volumes, it remains compatible with the isotropic formulation commonly used in prior work (2021; 2023; 2023; 2024; 2025).
+While cryoSplat adopts anisotropic Gaussians to represent 3D volumes, it remains compatible with the isotropic formulation commonly used in prior work (2021; 2023; 2023; 2024; 2025).
 As shown in Fig.

### v3

@@
 As a critical modality for structural biology, cryogenic electron microscopy (cryo-EM) facilitates the determination of macromolecular structures at near-atomic resolution.
-The core computational task in single-particle cryo-EM is to reconstruct the 3D electrostatic potential of a molecule from a large collection of noisy 2D projections acquired at unknown orientations.
+The core computational task in single-particle cryo-EM is to reconstruct the 3D electrostatic potential of a molecule from noisy 2D projections acquired at unknown orientations.
 Gaussian mixture models (GMMs) provide a continuous, compact, and physically interpretable representation for molecular density and have recently gained interest in cryo-EM reconstruction.
 However, existing methods rely on external consensus maps or atomic models for initialization, limiting their use in self-contained pipelines.
-Addressing this issue, we introduce cryoSplat, a GMM-based method that integrates Gaussian splatting with the physics of cryo-EM image formation.
-In particular, we develop an orthogonal projection-aware Gaussian splatting, with adaptations such as a normalization term and FFT-aligned coordinate system tailored for cryo-EM imaging.
-All these innovations enable stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images using random initialization.
+In parallel, differentiable rendering techniques such as Gaussian splatting have demonstrated remarkable scalability and efficiency for volumetric representations, suggesting a natural fit for GMM-based cryo-EM reconstruction.
+However, off-the-shelf Gaussian splatting methods are designed for photorealistic view synthesis and remain incompatible with cryo-EM due to mismatches in the image formation physics, reconstruction objectives, and coordinate systems.
+Addressing these issues, we propose cryoSplat, a GMM-based method that integrates Gaussian splatting with the physics of cryo-EM image formation.
+In particular, we develop an orthogonal projection-aware Gaussian splatting, with adaptations such as a view-dependent normalization term and FFT-aligned coordinate system tailored for cryo-EM imaging.
+These innovations enable stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images using random initialization.
 Experimental results on real datasets validate the effectiveness and robustness of cryoSplat over representative baselines.
 The code will be released upon publication.
-footnotetext: Preprint. Under review.
-Introduction
-Single particle cryogenic electron microscopy (cryo-EM) has emerged as a transformative tool in structural biology, enabling visualization of macromolecular complexes at atomic or near-atomic resolution without crystallization (Kühlbrandt 2014; Nogales 2016; Renaud et al.
-2018).
+1 Introduction
+Single particle cryogenic electron microscopy (cryo-EM) has emerged as a transformative tool in structural biology, enabling visualization of macromolecular complexes at atomic or near-atomic resolution without crystallization (Kühlbrandt, 2014; Nogales, 2016; Renaud et al., 2018).
 Central to cryo-EM is the computational task of reconstructing a 3D volume from a large collection of 2D projection images, each corresponding to a different, unknown viewing direction of identical particles embedded in vitreous ice.
-This inverse problem is inherently ill-posed and poses significant computational challenges.
+This inverse problem is inherently ill-posed and computationally challenging.
 First, cryo-EM images are severely corrupted by noise due to the low electron dose required to prevent radiation damage.
-For experimental datasets, the signal-to-noise (SNR) could be as low as around 20-20 dB (Bendory, Bartesaghi, and Singer 2020; Bepler et al.
-2020).
+For experimental datasets, the signal-to-noise (SNR) could be as low as around 20-20 dB (Bendory et al., 2020; Bepler et al., 2020).
 Second, the orientations (poses) of individual particles are unknown and must be inferred jointly with the 3D structure.
@@
 Both objectives fundamentally rely on the availability of a robust and efficient homogeneous reconstruction method, which assumes all particles correspond to a single structure and serves as a building block for more complex inference.
-Approaches to homogeneous reconstruction include methods based on backprojection, iterative expectation-maximization with voxel-based volumes (Tang et al.
-2007; Scheres 2012; Punjani et al.
-2017; Shekarforoush et al.
-2024), and more recently, neural representation learning (Zhong et al.
-2021a, b), which models the 3D volume using coordinate-based networks.
-In parallel, Gaussian mixture models (GMMs) have received attention for their continuous, compact, and physically interpretable parameterization of molecular density (Chen and Ludtke 2021; Chen, Schmid, and Chiu 2023).
-Notably, GMMs offer a natural connection to atomic models and can represent fine structural details using fewer parameters (Chen, Toader, and Roy 2023; Li et al.
-2024; Schwab et al.
-2024; Chen 2025).
-Despite their conceptual appeal, existing GMM-based methods for cryo-EM reconstruction rely on nontrivial prerequisite steps.
-These approaches (2021; 2023; 2023; 2024; 2024; 2025) typically rely on consensus volumes from external pipelines, or even atomic models, for initialization, and have not demonstrated stable convergence when directly optimizing from experimental images.
+Approaches to homogeneous reconstruction include methods based on backprojection, iterative expectation-maximization with voxel-based volumes (Tang et al., 2007; Scheres, 2012; Punjani et al., 2017; Shekarforoush et al., 2024), and more recently, neural representation learning (Zhong et al., 2021a; b), which models the 3D volume using coordinate-based networks.
+In parallel, Gaussian mixture models (GMMs) have received attention for their continuous, compact, and physically interpretable parameterization of molecular density (Chen & Ludtke, 2021; Chen et al., 2023a).
+Notably, GMMs offer a natural connection to atomic models and can represent fine structural details using fewer parameters (Chen et al., 2023b; Li et al., 2024; Schwab et al., 2024; Chen, 2025).
+Despite their conceptual appeal, existing GMM-based methods (Chen & Ludtke, 2021; Chen et al., 2023a; b; Li et al., 2024; Schwab et al., 2024; Chen, 2025) for cryo-EM reconstruction rely on nontrivial prerequisite steps.
+These approaches typically rely on consensus volumes from external pipelines, or even atomic models, for initialization, and have not demonstrated stable convergence when directly optimizing from experimental images.
 In fact, no prior method achieves reliable GMM-based reconstruction even under known particle poses, due to the inherent difficulty of optimizing mixture parameters in extreme noise.
@@
 Given known particle poses, cryoSplat performs stable and efficient reconstruction directly from raw cryo-EM projection images, starting from random initialization and requiring no external priors.
-Inspired by recent advances in 3D Gaussian Splatting (3DGS) (Kerbl et al.
-2023), we model the 3D density as a mixture of anisotropic Gaussians and project them into 2D using a novel differentiable orthographic splatting algorithm consistent with cryo-EM physics.
+Inspired by recent advances in 3D Gaussian Splatting (3DGS) by Kerbl et al.
+(2023), we model the 3D density as a mixture of anisotropic Gaussians and project them into 2D using a novel differentiable orthographic splatting algorithm consistent with cryo-EM physics.
 To support practical scalability and training efficiency, we develop a CUDA-accelerated real-space renderer that enables fast rasterization and optimization of the GMM.
 Our contributions can be summarized as follows:
-  • A self-contained GMM-based reconstruction method: We present cryoSplat as the first method capable of performing cryo-EM homogeneous reconstruction from scratch using only a Gaussian Mixture Model, thereby establishing the missing foundation needed to develop GMMs into standalone reconstruction tools.
+  • A self-contained GMM-based reconstruction method: We present cryoSplat as the first method capable of performing cryo-EM homogeneous reconstruction from a randomly initialized Gaussian mixture model without an external prior, thereby establishing the missing foundation needed to develop GMMs into standalone reconstruction tools.
   • A physically accurate projection model: We design a splatting algorithm under orthogonal projection tailored to cryo-EM image formation, enabling differentiable projection of anisotropic Gaussians in real space.
@@
   • Experimental validation: We demonstrate the effectiveness of cryoSplat on real datasets, showing it converges reliably from random initialization and achieves reconstruction quality outperforming state-of-the-art methods.
-Related Work
-Volume Representation in Cryo-EM
+2 Related work
+2.1 Volume representation in cryo-EM
 In cryo-EM experiments, purified biomolecules are rapidly frozen in a thin layer of vitreous ice, where each particle adopts a random orientation.
-A high-energy electron beam passes through the specimen, interacts with the electrostatic potential of the particles, and is recorded on a detector as a 2D projection image (Singer and Sigworth 2020).
+A high-energy electron beam passes through the specimen, interacts with the electrostatic potential of the particles, and is recorded on a detector as a 2D projection image (Singer & Sigworth, 2020).
 The goal of cryo-EM reconstruction is to recover the 3D electrostatic potential, i.e., the volume, from a large set of such noisy and randomly oriented 2D projections.
 Central to cryo-EM reconstruction is the choice of volume representation.
-Voxel-based representations
-are the most widely used in conventional cryo-EM software, e.g., RELION (2016), cryoSPARC (2017) and EMAN2 (2007).
+2.1.1 Voxel-based representation
+Voxel-based representations are the most widely used in conventional cryo-EM software, e.g., RELION (Scheres, 2016b), cryoSPARC (Punjani et al., 2017) and EMAN2 (Tang et al., 2007).
 The 3D volume is discretized into a regular grid of density values, enabling fast projection and reconstruction via FFT-based algorithms.
 Despite their practical success, voxel grids are memory-intensive and inherently discrete, which limits their compatibility with modern learning-based analysis frameworks.
-Neural fields
-represent the volume as a continuous function parameterized by neural networks.
-These methods (Zhong et al.
-2021a, b; Levy et al.
-2022a, b, 2025) offer differentiability, implicit smoothness, and natural compatibility with learning-based heterogeneous analysis.
+2.1.2 Neural field
+Neural fields represent the volume as a continuous function parameterized by neural networks.
+These methods (Zhong et al., 2021a; b; Levy et al., 2022a; b; 2025) offer differentiability, implicit smoothness, and natural compatibility with learning-based heterogeneous analysis.
 However, the implicit nature of neural fields often comes at the cost of interpretability, and such models are typically slow to train and difficult to constrain with biological priors.
-Gaussian mixture models
-have a long history in structural biology, with early uses for molecular approximation (Grant and Pickup 1995; Grant, Gallardo, and Pickup 1996; Kawabata 2008).
-E2GMM (Chen and Ludtke 2021) was among the first to apply GMMs to cryo-EM heterogeneous reconstruction.
+2.1.3 Gaussian mixture model
+Gaussian mixture models have a long history in structural biology, with early uses for molecular approximation (Grant & Pickup, 1995; Grant et al., 1996; Kawabata, 2008).
+E2GMM (Chen & Ludtke, 2021) was among the first to apply GMMs to cryo-EM heterogeneous reconstruction.
 Like neural fields, GMMs can approximate any smooth density function and support differentiable optimization.
 More importantly, GMMs provide an explicit and interpretable representation that naturally links to atomic structures.
-Recent studies (Chen, Schmid, and Chiu 2023; Chen, Toader, and Roy 2023; Li et al.
-2024; Schwab et al.
-2024; Chen 2025) have shown that GMMs can capture molecular flexibility by modeling atomic motion directly, making them highly suitable for heterogeneous reconstruction and downstream structural analysis.
+Recent studies (Chen et al., 2023a; b; Li et al., 2024; Schwab et al., 2024; Chen, 2025) have shown that GMMs can capture molecular flexibility by modeling atomic motion directly, making them highly suitable for heterogeneous reconstruction and downstream structural analysis.
 However, existing GMM-based methods typically require initialization from an externally reconstructed consensus map or even an atomic model.
@@
 Our work addresses this limitation by introducing a GMM-based reconstruction pipeline that can be stably trained from scratch.
-Gaussian Splatting
-3DGS (Kerbl et al.
-2023) is a recent differentiable rendering technique developed for real-time novel view synthesis.
-It represents a 3D scene as a collection of anisotropic Gaussians and renders images via rasterization-based accumulation and alpha blending (Zwicker et al.
-2002).
-While 3DGS achieves high visual fidelity in synthetic and real-world RGB datasets, as a volume rendering method, it is not a physically accurate model of natural image formation (Huang et al.
-2024).
+2.2 Gaussian splatting
+3DGS (Kerbl et al., 2023) is a recent differentiable rendering technique developed for real-time novel view synthesis.
+It represents a 3D scene as a collection of anisotropic Gaussians and renders images via rasterization-based accumulation and alpha blending (Zwicker et al., 2002).
+While 3DGS achieves high visual fidelity in synthetic and real-world RGB datasets, as a volume rendering method, it is not a physically accurate model of natural image formation (Huang et al., 2024).
 Although the original 3DGS formulation is not physically consistent with natural image formation, its volume rendering framework closely aligns with the cryo-EM imaging model, where each image arises from an orthographic line integral of electrostatic potential modulated by the contrast transfer function (CTF).
 Leveraging this alignment, we adapt splatting to cryo-EM by rederiving the projection of anisotropic Gaussians under cryo-EM physics, replacing heuristic alpha blending with physically accurate line integrals and incorporating CTF modulation.
-Method
+3 Method
 Figure 1: Cryo-EM reconstruction aims to recover a 3D volume (a) from a large set of 2D particle images (b).
@@
 The resulting GMM can be voxelized to obtain the final 3D volume.
-Overview
+3.1 Overview
 Our goal is to achieve physically accurate and computationally efficient cryo-EM reconstruction by leveraging a Gaussian Mixture Model (GMM).
 To this end, we propose cryoSplat, a differentiable framework that represents the 3D electrostatic potential of a specimen as a set of anisotropic Gaussians and directly simulates the cryo-EM image formation process in real space, faithfully adhering to the physics of transmission electron microscopy.
-Building upon recent progress in differentiable volume rendering, particularly the Gaussian splatting framework introduced in 3DGS (2023), we adopt a tile-based rasterization strategy to enable scalable and efficient computation.
+Building upon recent progress in differentiable volume rendering, particularly the Gaussian splatting framework by Kerbl et al.
+(2023), we adopt a tile-based rasterization strategy for scalable and efficient computation.
 However, the original 3DGS formulation is not directly applicable to cryo-EM due to several fundamental mismatches: (i) it employs perspective projection consistent with a pinhole camera model, in contrast to the orthographic projection in cryo-EM imaging; (ii) it prioritizes photorealistic novel view synthesis over physical reconstruction accuracy; and (iii) its coordinate system is misaligned with the Fourier slice theorem, which underpins cryo-EM reconstruction.
-To address these issues, cryoSplat introduces several key adaptations: (i) we replace the heuristic alpha compositing with physically grounded line integrals to reflect the transmission nature of electron imaging; (ii) we fix the normalization between 3D-to-2D transformation and apply consistent learning rates across all parameters to ensure stable optimization; and (iii) we align the rasterization coordinate system with the FFT grid, allowing accurate gradient propagation through CTF modulation.
-These modifications collectively enable cryoSplat to perform stable, end-to-end differentiable reconstruction from raw cryo-EM particle images, starting from random initialization without relying on externally provided volumes.
-Image Formation
+To address these issues, cryoSplat introduces several key adaptations: (i) we replace heuristic alpha blending with physically grounded line integrals to reflect the transmission nature of electron imaging; (ii) we fix the normalization between 3D-to-2D transformation and apply consistent learning rates across all parameters to ensure stable optimization; and (iii) we align the rasterization coordinate system with the FFT grid, allowing accurate gradient propagation through CTF modulation.
+These modifications collectively enable cryoSplat to perform stable, end-to-end differentiable reconstruction from raw cryo-EM particle images, starting from random initialization without relying on externally provided volumes or atomic models.
+3.2 Image formation
 As shown in Fig.
-1(b), a beam of electrons passes through a vitrified specimen, and the transmitted wavefront undergoes phase shifts due to the specimen’s electrostatic potential(Singer and Sigworth 2020).
-Under the weak phase approximation, these phase shifts are linearly related to the 3D potential (volume), and the image formed at the detector is a line integral (projection) of this potential along the beam direction, further convolved with $\mathcal{H}$, the point spread function (PSF) of the imaging system.
-In homogeneous reconstruction, it is assumed that all particle images $\mathcal{I}$ correspond to identical copies of a single 3D volume $\mathcal{V}:\mathbb{R}^{3}\rightarrow\mathbb{R}$, and that any conformational or compositional heterogeneity is negligible.
+1(b), electrons traverse a vitrified specimen, and the transmitted wavefronts undergo phase shifts due to the specimen’s electrostatic potential (Singer & Sigworth, 2020).
+Under the weak phase approximation, the phase shifts are linearly related to the 3D potential (volume), and the image formed at the detector is a line integral (projection) of this potential along the beam direction, further convolved with $H\!:{\mathbb{R}}^{2}\!\rightarrow\!{\mathbb{R}}$, the point spread function (PSF) of the imaging system.
+In homogeneous reconstruction, it is assumed that all particle images $Y:{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ correspond to identical copies of a single 3D volume $V:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}$, and that any conformational or compositional heterogeneity is negligible.
 Under this assumption, the image formation model can be expressed as:
-$\mathcal{I}(r_{x},r_{y})=\mathcal{H}(r_{x},r_{y})*\int_{\mathbb{R}}\mathcal{V}(\mathbf{W}^{\top}\mathbf{r}+\mathbf{t})dr_{z}+\epsilon,$ (1)
-where $\mathbf{r}=[r_{x},r_{y},r_{z}]^{\top}$ denotes the 3D Cartesian coordinates in real space, $\mathbf{W}\in SO(3)$ is the 3D pose of the volume, and $\mathbf{t}=[t_{x},t_{y},0]^{\top}$ is the in-plane translation, accounting for imperfect centering during particle cropping.
-The noise term $\epsilon$ is typically modeled as independent, zero-centered Gaussian noise.
-According to the Fourier slice theorem (Bracewell 1956), illustrated in Fig.
-1(c), the 2D Fourier transform of a projection corresponds to a central slice of the 3D Fourier transform of the volume, orthogonal to the projection direction and passing through the origin.
-Based on this property, an alternative and widely adopted formulation models reconstruction directly in the Fourier domain, where the image formation model becomes:
-$\widehat{\mathcal{I}}(k_{x},k_{y})=\widehat{\mathcal{H}}(k_{x},k_{y})\cdot\widehat{\mathcal{V}}(\mathbf{W}^{\top}\mathbf{k})\cdot e^{-2\pi i\mathbf{k}^{\top}\mathbf{t}}+\widehat{\epsilon}$, (2)
-where $\mathbf{k}=[k_{x},k_{y},0]^{\top}$ denotes the Cartesian coordinates in Fourier space, and the 2D spectrum $\widehat{\mathcal{I}}$, the CTF $\widehat{\mathcal{H}}$ and the 3D spectrum $\widehat{\mathcal{V}}$ denote the Fourier transform of $\mathcal{I}$, $\mathcal{H}$ and $\mathcal{V}$, respectively.
-The noise term $\widehat{\epsilon}$ is similarly modeled as independent, zero-mean Gaussian noise in the Fourier domain.
-In this work, departing from most existing approaches that adopt Eq. (2), we instead build our pipeline on Eq. (1), performing homogeneous reconstruction directly in real space.
-CryoSplat
-Anisotropic GMMs
-are developed to represent the volume, which can be written in the form
-$\mathcal{V}(\mathbf{r})=\sum_{i=1}^{N}A_{i}\mathcal{G}_{i}(\mathbf{r}),$ (3)
-where $N$ denotes the number of Gaussians and $A_{i}$ is the amplitude of the $i$-th normalized Gaussian $\mathcal{G}_{i}(\mathbf{r})$.
+$Y(r_{x},r_{y})=H(r_{x},r_{y})*\int_{{\mathbb{R}}}V({\bm{W}}^{\top}{\bm{r}}+{\bm{t}})dr_{z}+{\textnormal{$\epsilon$}},$ (1)
+where ${\bm{r}}=[r_{x},r_{y},r_{z}]^{\top}$ are the 3D Cartesian coordinates in real space, ${\bm{W}}\in\mathrm{SO}(3)$ is the 3D pose of the particle, and ${\bm{t}}=[t_{x},t_{y},0]^{\top}$ is the in-plane translation, accounting for imperfect centering during particle cropping.
+The noise term $\epsilon$ is modeled as independent, zero-mean Gaussian noise.
+3.3 CryoSplat
+3.3.1 Anisotropic GMM
+Anisotropic GMMs are developed to represent the volume, which can be written in the form
+$V({\bm{r}})=\sum_{i=1}^{N}A_{i}G_{i}({\bm{r}}),$ (2)
+where $N$ denotes the Guassian count and $A_{i}$ is the amplitude of the $i$-th normalized Gaussian $G_{i}({\bm{r}})$.
 By substituting Eq.
@@
 Specifically, we apply a viewing transformation to align the GMM to the target orientation, orthographically project each Gaussian along the $z$-axis to form a 2D image, and convolve the result with the PSF:
-$\mathcal{I}(r_{x},r_{y})=\mathcal{H}(r_{x},r_{y})*\sum_{i=1}^{N}A_{i}\int_{\mathbb{R}}\mathcal{G}_{i}(\mathbf{W}^{\top}\mathbf{r}+\mathbf{t})dr_{z}.$ (4)
+$X(r_{x},r_{y})=H(r_{x},r_{y})*\sum_{i=1}^{N}A_{i}\int_{{\mathbb{R}}}G_{i}({\bm{W}}^{\top}{\bm{r}}+{\bm{t}})dr_{z}.$ (3)
 Since the integral is linear, each Gaussian contributes independently to the final image.
@@
 A normalized 3D Gaussian is defined as:
-$\mathcal{G}(\mathbf{r}|\boldsymbol{\mu},\boldsymbol{\Sigma})=\frac{1}{(2\pi)^{\frac{3}{2}}|\boldsymbol{\Sigma}|^{\frac{1}{2}}}\exp(-\frac{1}{2}(\mathbf{r}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}(\mathbf{r}-\boldsymbol{\mu})),$ (5)
-where $\boldsymbol{\mu}\in\mathbb{R}^{3}$ and $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$ denote the mean (center) and the covariance matrix, respectively.
-The determinant $|\boldsymbol{\Sigma}|$ ensures proper normalization.
-Following (Kerbl et al.
-2023), to guarantee the positive semidefinite property, we construct the covariance matrix as:
-$\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top},$ (6)
-where $\mathbf{S}=\mathrm{diag}(\mathbf{s})$ is a diagonal scaling matrix and $\mathbf{R}\in SO(3)$ is a rotation matrix.
-In our implementation, we store the scaling vector $\mathbf{s}=[s_{x},s_{y},s_{z}]^{\top}$ and parameterize $\mathbf{R}$ using a quaternion $\mathbf{q}=[q_{w},q_{x},q_{y},q_{z}]^{\top}$.
-To ensure positivity and stable gradients during optimization, we apply a softplus function to both the amplitude $A$ and the scaling vector $\mathbf{s}$.
-The quaternion $\mathbf{q}$ is normalized to ensure it represents a valid rotation.
+$G({\bm{r}}|{\bm{\mu}},{\bm{\Sigma}})=\frac{1}{(2\pi)^{\frac{3}{2}}|{\bm{\Sigma}}|^{\frac{1}{2}}}\exp\Big(-\frac{1}{2}({\bm{r}}-{\bm{\mu}})^{\top}{\bm{\Sigma}}^{-1}({\bm{r}}-{\bm{\mu}})\Big),$ (4)
+where ${\bm{\mu}}\in{\mathbb{R}}^{3}$ and ${\bm{\Sigma}}\in{\mathbb{R}}^{3\times 3}$ denote the mean (position) and the covariance matrix (shape), respectively.
+The determinant $|{\bm{\Sigma}}|$ ensures proper normalization.
+Following Kerbl et al.
+(2023), to guarantee the positive semidefinite property, we construct the covariance matrix as:
+${\bm{\Sigma}}={\bm{R}}{\bm{S}}{\bm{S}}^{\top}{\bm{R}}^{\top},$ (5)
+where ${\bm{S}}=\mathrm{diag}({\bm{s}})$ is a diagonal scaling matrix and ${\bm{R}}\in\mathrm{SO}(3)$ is a rotation matrix.
+In our implementation, we store the scaling vector ${\bm{s}}=[s_{x},s_{y},s_{z}]^{\top}$ and parameterize $\bm{R}$} using a quaternion ${\bm{q}}=[q_{w},q_{x},q_{y},q_{z}]^{\top}$.
+To ensure positivity and stable gradients during optimization, we apply a softplus function to both the amplitude $A$ and the scaling vector $\bm{s}$}.
+The quaternion $\bm{q}$} is normalized to ensure it represents a valid rotation.
 Altogether, each anisotropic Gaussian is parameterized by the 11-dimensional set $\{\mu_{x},\mu_{y},\mu_{z},s_{x},s_{y},s_{z},q_{w},q_{x},q_{y},q_{z},A\}$.
-When the scaling is isotropic, i.e., $s_{x}=s_{y}=s_{z}=\sigma$, the anisotropic Gaussian reduces to the standard isotropic form:
-$\mathcal{G}(\mathbf{r}|\boldsymbol{\mu},\sigma)=\frac{1}{(2\pi)^{\frac{3}{2}}\sigma}\exp(-\frac{\|\mathbf{r}-\boldsymbol{\mu}\|_{2}^{2}}{2\sigma^{2}}),$ (7)
-which is more commonly used in previous cryo-EM reconstruction works (2021; 2023; 2023; 2024; 2024; 2025).
-The viewing transformation
-is the first step in simulating image formation, aligning each Gaussian with a given projection direction.
-Since the parameters $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$ describe Gaussians in world coordinates, we must transform them into the image-space coordinates before projection.
-According to the derivation in (Zwicker et al.
-2002), applying an affine transformation to a Gaussian results in another Gaussian with appropriately transformed parameters.
-In our case, the transformation consists of a rotation $\mathbf{W}\in SO(3)$ and a 2D in-plane translation $\mathbf{t}\in\mathbb{R}^{3}$, leading to:
-$\dot{\mathcal{G}}(\mathbf{r}|\dot{\boldsymbol{\mu}},\dot{\boldsymbol{\Sigma}})=\mathcal{G}(\mathbf{W}^{\top}\mathbf{r}+\mathbf{t}|\boldsymbol{\mu},\boldsymbol{\Sigma}),$ (8)
-where the transformed mean and covariance are given by $\dot{\boldsymbol{\mu}}=\mathbf{W}\boldsymbol{\mu}+\mathbf{t}$ and $\dot{\boldsymbol{\Sigma}}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}$, respectively.
-The orthogonal projection
+3.3.2 Viewing transformation
+The viewing transformation is the first step in simulating image formation, aligning each Gaussian with a given projection direction.
+Since the parameters $\bm{\mu}$} and $\bm{\Sigma}$} describe Gaussians in world coordinates, we must transform them into the image-space coordinates before projection.
+According to the derivation in Zwicker et al.
+(2002), applying an affine transformation to a Gaussian results in another Gaussian with appropriately transformed parameters.
+In our case, the transformation consists of a rotation ${\bm{W}}\in\mathrm{SO}(3)$ and a 2D in-plane translation ${\bm{t}}\in{\mathbb{R}}^{3}$, leading to:
+$\dot{G}({\bm{r}}|\dot{{\bm{\mu}}},\dot{{\bm{\Sigma}}})=G({\bm{W}}^{\top}{\bm{r}}+{\bm{t}}|{\bm{\mu}},{\bm{\Sigma}}),$ (6)
+where the transformed mean and covariance are given by $\dot{{\bm{\mu}}}={\bm{W}}({\bm{\mu}}-{\bm{t}})$ and $\dot{{\bm{\Sigma}}}={\bm{W}}{\bm{\Sigma}}{\bm{W}}^{\top}$.
+3.3.3 Orthogonal projection
 […]
-$\tilde{\mathcal{G}}(\tilde{\mathbf{r}}|\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}})=\int_{\mathbb{R}}\dot{\mathcal{G}}(\mathbf{r}|\dot{\boldsymbol{\mu}},\dot{\boldsymbol{\Sigma}})dr_{z}.$ (9)
+$\tilde{G}(\tilde{{\bm{r}}}|\tilde{{\bm{\mu}}},\tilde{{\bm{\Sigma}}})=\int_{{\mathbb{R}}}\dot{G}({\bm{r}}|\dot{{\bm{\mu}}},\dot{{\bm{\Sigma}}})dr_{z}.$ (7)
 This operation effectively integrates the 3D Gaussian along the projection axis, preserving its Gaussian form in 2D. The resulting closed-form expression is:
-$),\tilde{\mathcal{G}}(\tilde{\mathbf{r}}|\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}})=\frac{1}{2\pi|\tilde{\boldsymbol{\Sigma}}|^{\frac{1}{2}}}\exp(-\frac{1}{2}(\tilde{\mathbf{r}}-\tilde{\boldsymbol{\mu}})^{\top}\tilde{\boldsymbol{\Sigma}}^{-1}(\tilde{\mathbf{r}}-\tilde{\boldsymbol{\mu}}$)), (10)
-where $\tilde{\mathbf{r}}=[r_{x},r_{y}]^{\top}$ denotes the 2D Cartesian coordinates in real space.
-As discussed in (Zwicker et al.
-2002), this projection can be computed analytically by removing the z-axis components from the mean and covariance:
-$\begin{cases}\dot{\boldsymbol{\mu}}=[\dot{\mu}_{x},\dot{\mu}_{y},\dot{\mu}_{z}]^{\top}\Rightarrow[\dot{\mu}_{x},\dot{\mu}_{y}]^{\top}=\tilde{\boldsymbol{\mu}}\\ \dot{\boldsymbol{\Sigma}}=\begin{bmatrix}\dot{\sigma}_{xx}&\dot{\sigma}_{xy}&\dot{\sigma}_{xz}\\ \dot{\sigma}_{xy}&\dot{\sigma}_{yy}&\dot{\sigma}_{yz}\\ \dot{\sigma}_{xz}&\dot{\sigma}_{yz}&\dot{\sigma}_{zz}\\ \end{bmatrix}\Rightarrow\begin{bmatrix}\dot{\sigma}_{xx}&\dot{\sigma}_{xy}\\ \dot{\sigma}_{xy}&\dot{\sigma}_{yy}\\ \end{bmatrix}=\tilde{\boldsymbol{\Sigma}}\end{cases}$ (11)
-In prior works, such as 3DGS, the normalization term $1/(2\pi|\tilde{\boldsymbol{\Sigma}}|^{\frac{1}{2}})$ is often omitted, as their primary focus is on photorealistic novel view synthesis rather than the physical fidelity of the underlying 3D representation.
+$),\tilde{G}(\tilde{{\bm{r}}}|\tilde{{\bm{\mu}}},\tilde{{\bm{\Sigma}}})=\frac{1}{2\pi|\tilde{{\bm{\Sigma}}}|^{\frac{1}{2}}}\exp\Big(-\frac{1}{2}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})^{\top}\tilde{{\bm{\Sigma}}}^{-1}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}$$\Big)$, (8)
+where $\tilde{{\bm{r}}}=[r_{x},r_{y}]^{\top}$ denotes the 2D Cartesian coordinates in real space.
+In prior works, such as 3DGS, the normalization term $1/(2\pi|\tilde{{\bm{\Sigma}}}|^{\frac{1}{2}})$ is often omitted, as their primary focus is on photorealistic novel view synthesis rather than the physical fidelity of the underlying 3D representation.
 However, in cryo-EM reconstruction, the ultimate goal is to recover the correct 3D volume.
-Omitting this normalization introduces bias in amplitude and leads to incorrect reconstructions.
+Omitting this view-dependent normalization introduces bias in amplitude and leads to incorrect reconstructions.
 Therefore, unlike 3DGS, we retain the normalization term to preserve the quantitative correctness of the model.
-After projection, the final image is constructed by summing the weighted contributions of all projected 2D Gaussians and applying the PSF:
+After projection, the final image is constructed by summing the weighted contributions of all splats and applying the PSF:
 […]
-In practice, consistent with typical cryo-EM pipelines, we apply the CTF instead in the Fourier domain:
-[…]
-where $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the Fourier and inverse Fourier transform operators, respectively.
-Fast differentiable rasterization.
-We adopt the efficient tile-based rasterization framework from 3DGS (Kerbl et al.
-2023), which enables scalable and differentiable processing of tens of thousands of Gaussians via per-tile accumulation.
-Unlike 3DGS, which uses alpha blending for photorealistic rendering, we modify the rasterization to directly sum projected Gaussian contributions, in accordance with the physical transmission model in cryo-EM.
-A key reason we choose to operate in real space is that it allows us to fully exploit this fast rasterization strategy.
-In high-resolution reconstructions, individual Gaussians in real space have small spatial scales and affect only a few nearby pixels.
-This locality means each GPU thread—responsible for a single pixel—only needs to process a small subset of all Gaussians.
-In contrast, Gaussians in Fourier space become broad as resolution increases, leading to near-global support.
-As a result, each pixel in the frequency domain must aggregate contributions from nearly all Gaussians, making fast rendering impractical.
-For an image of shape $D\times D$, the original 3DGS implementation places the continuous coordinate center at $(D-1)/2,(D-1)/2]^{\top}$, i.e., halfway between two discrete pixels.
+3.3.4 Fast differentiable rasterization
+We adopt the efficient tile-based rasterization framework from Kerbl et al.
+(2023), which enables scalable and differentiable processing of tens of thousands of Gaussians via per-tile accumulation.
+Unlike 3DGS, which uses alpha blending for photorealistic rendering, we modify the rasterization to directly sum contributions of splats, in accordance with the physical transmission model in cryo-EM.
+For an image ${\bm{X}}\!\in\!{\mathbb{R}}^{D\times D}$, the original 3DGS implementation places the continuous coordinate center at $(D-1)/2,(D-1)/2]^{\top}$, i.e., halfway between two discrete pixels.
 In contrast, FFT-based image formation assumes the origin is located at the integer grid point $\lfloor D/2\rfloor,\lfloor D/2\rfloor]^{\top}$.
@@
 This alignment eliminates phase inconsistencies and enables accurate electron projection simulation, while preserving the computational efficiency of the 3DGS architecture.
-Loss function.
-Unlike previous GMM-based methods, which often rely on specially designed loss functions combined with complex regularization or constraints to ensure stable optimization, we adopt a much simpler formulation.
-Specifically, we directly apply the mean squared error (MSE) loss between the GMM-based projection $\mathcal{I}$ and the observed image $\mathcal{I}_{\rm ref}$:
-$\mathcal{L}=\|\mathcal{I}-\mathcal{I}_{\rm ref}\|_{2}.$ (14)
+Let ${\bm{X}},{\bm{Y}}\!\in\!\mathbb{R}^{D\times D}$ be the matrices representing the GMM-based projection $X$ and the observed image $Y$ after rasterization, respectively.
+3.3.5 Loss function
+Unlike previous GMM-based methods that rely on specially designed losses with complex regularization or constraints to ensure stable optimization, we adopt a much simpler formulation.
+Specifically, we directly apply the mean squared error (MSE) loss between the GMM-based projection $\bm{X}$} and the observed image $\bm{Y}$}: $\mathcal{L}=\|{\bm{X}}-{\bm{Y}}\|_{2}^{2}$.
 Despite its simplicity, this loss formulation leads to stable and fast convergence in practice, without requiring additional regularization terms.
-Experiment
+4 Experiment
 Figure 2: Qualitative and quantitative comparison of voxel-based, neural, and GMM-based representations.
-(Left) Final 3D reconstructions on four real datasets visualized with ChimeraX (Pettersen et al.
-2021).
+(Left) Final 3D reconstructions on four real datasets visualized with ChimeraX (Pettersen et al., 2021).
 (Right) FSC curves are plotted for quantitative evaluation.
@@
 CryoSplat consistently achieves higher resolution across all datasets.
-Figure 3: Evaluation of reconstruction performance with varying numbers of Gaussians.
-Increasing the number improves both accuracy and robustness, with 10,00010{,}000 Gaussians generally sufficient for most datasets.
-Figure 4: Runtime efficiency comparison across reconstruction methods at different resolutions ($D=192$ and $D=256$). Frame rates (FPS) are measured under increasing numbers of Gaussians (log-scaled).
-Experimental Settings
+Figure 3: Reconstruction performance with varying numbers of Gaussians. Increasing the number improves accuracy and robustness, with 10,00010{,}000 Gaussians generally sufficient for most datasets.
+Figure 4: Runtime efficiency across reconstruction methods at different resolutions ($D=192$ and $D=256$). Frame rates (FPS) are measured under increasing numbers of Gaussians (log-scaled).
+4.1 Experimental settings
 Datasets.
-We evaluate our method on four publicly available cryo-EM datasets from the Electron Microscopy Public Image Archive (EMPIAR) (Iudin et al.
-2016): EMPIAR-10028 (Pf80S ribosome) (Wong et al.
-2014), EMPIAR-10049 (RAG complex) (Ru et al.
-2015), EMPIAR-10076 (E.
-coli LSU assembly) (Davis et al.
-2016), and EMPIAR-10180 (pre-catalytic spliceosome) (Plaschka, Lin, and Nagai 2017).
+We evaluate our method on four publicly available cryo-EM datasets from the Electron Microscopy Public Image Archive (EMPIAR) (Iudin et al., 2016): EMPIAR-10028 (Pf80S ribosome) (Wong et al., 2014), EMPIAR-10049 (RAG complex) (Ru et al., 2015), EMPIAR-10076 (E.
+coli LSU assembly) (Davis et al., 2016), and EMPIAR-10180 (pre-catalytic spliceosome) (Plaschka et al., 2017).
 These datasets span a range of structural complexity and image quality, from rigid assemblies with high contrast to highly heterogeneous macromolecular machines.
@@
 Evaluation metrics.
-Since ground truth volumes are unavailable for real datasets, we follow standard practice and assess reconstruction quality using the gold standard Fourier Shell Correlation (FSC) (Van Heel and Schatz 2005).
-Each dataset is split evenly into two halves, and the method is applied independently to each subset.
-The resulting reconstructions are compared in Fourier space by computing FSC as a function of spatial frequency, which quantifies their consistency across frequency shells.
-The resolution is defined as the spatial frequency at which the FSC curve drops below the $0.143$ threshold.
+Since ground truth volumes are unavailable for real datasets, we follow standard practice and assess reconstruction quality using the gold standard Fourier Shell Correlation (FSC) (Van Heel & Schatz, 2005).
+Each dataset is split evenly into two halves, and the method is applied independently to each.
+Reconstructed volumes are compared in Fourier space by computing FSC as a function of spatial frequency, quantifying their consistency across frequency shells.
+Resolution is defined as the spatial frequency where the FSC curve drops below the $0.143$ threshold.
 Implementation details.
@@
 Published particle translations are applied to the observed images via phase shifting in Fourier space prior to reconstruction, rather than through the GMM viewing transform.
-The 3D volume domain is defined as $0.5,0.5]^{3}$, assuming each projection lies within $0.5,0.5]^{2}$ in the image plane.
-Gaussians are initialized with random means $\boldsymbol{\mu}\sim\mathcal{N}(\mathbf{0},0.075^{2}\mathbf{I})$, isotropic scales $s_{x}\!=\!s_{y}\!=\!s_{z}\!=\!0.0075$, identity quaternion $\mathbf{q}\!=\![1,0,0,0]^{T}$, and amplitude $A=1/(2N)$, where $N$ is the number of Gaussians.
+We do not apply any windowing to the observed particle images during preprocessing.
+The 3D volume is defined over the domain $E,E]^{3}$, and each 2D projection is assumed to lie within $E,E]^{2}$ in the image plane, where $E=0.5$ defines the spatial extent.
+Gaussians are initialized with random means ${\bm{\mu}}\sim\mathcal{N}({\bm{0}},0.075^{2}{\bm{I}})$, isotropic scales $s_{x}\!=\!s_{y}\!=\!s_{z}\!=\!0.0075$, identity quaternion ${\bm{q}}\!=\![1,0,0,0]^{T}$, and amplitude $A=1/(2N)$, where $N$ is the number of Gaussians.
 All parameters are trainable.
-We use Adam (Kingma and Ba 2014) with batch size $1$, learning rate $0.001$, and exponential decay ($\gamma=0.1$) at each epoch.
+We use Adam (Kingma & Ba, 2014) with batch size $1$, learning rate $0.001$, and exponential decay ($\gamma=0.1$) at each epoch.
 All GMMs are trained for $5$ epochs.
-For the voxel-based backprojection baseline, we use cryoDRGN’s official implementation (Zhong et al.
-2021a).
+For the voxel-based backprojection baseline, we use the implementation provided by Zhong et al.
+(2021a).
 For neural representation learning, we follow cryoDRGN’s default configuration: three 1,0241{,}024-node layers, trained for $50$ epochs.
-Since related methods (2021a; 2021b; 2022a; 2022b; 2025) share similar neural fields and differ mainly in pose estimation, we focus comparison on cryoDRGN.
 All experiments are run on a single NVIDIA GeForce RTX 3090.
-Evaluation on Real Datasets
+4.2 Evaluation on real datasets
 We evaluate the performance of different volume representations on real cryo-EM datasets under a homogeneous reconstruction setting.
@@
 Our evaluation focuses on two aspects: (i) the ability to reconstruct high-resolution consensus maps, and (ii) robustness to noise and imperfect pose assignments.
-We compare voxel-based backprojection, the neural representation method cryoDRGN (Zhong et al.
-2021a), and our proposed GMM-based method cryoSplat.
+Since related methods (Zhong et al., 2021a; b; Levy et al., 2022a; b; 2025) adopt cryoDRGN’s neural field implementation and differ mainly in pose estimation, we focus our comparison on cryoDRGN.
+Accordingly, we evaluate three representative approaches: voxel-based backprojection, the neural representation method cryoDRGN (Zhong et al., 2021a), and our proposed GMM-based method cryoSplat.
 Visualizations of the reconstructed volumes are shown in Fig.
@@
 Each volume of cryoSplat is represented using 30,00030{,}000 Gaussians.
-EMPIAR-10028.
-The Pf80S ribosome is relatively easy to reconstruct due to its high-contrast images and structurally stable particles.
+The Pf80S ribosome (EMPIAR-10028) is relatively easy to reconstruct due to its high-contrast images and structurally stable particles.
 All methods achieve high-resolution results ($3.80$ Å) and strong FSC agreement across the spectrum.
 cryoDRGN yields slightly higher FSC values at intermediate frequencies, while cryoSplat outperforms all baselines at high spatial frequencies, demonstrating its ability to recover fine structural details.
-EMPIAR-10049.
-The RAG complex poses greater challenges due to imperfect published poses and flexible regions such as the DNA elements and the nonamer binding domain (NBD), indicated by arrows.
+The RAG complex (EMPIAR-10049) poses greater challenges due to imperfect published poses and flexible regions such as the DNA elements and the nonamer binding domain (NBD), indicated by arrows.
 CryoSplat significantly outperforms both backprojection and cryoDRGN, achieving a resolution of $2.49$ Å.
@@
 Its FSC curve remains consistently above those of other methods across all spatial frequencies, highlighting its robustness to pose inaccuracies and structural variability.
-EMPIAR-10076.
-This LSU assembly dataset contains substantial compositional and conformational heterogeneity, making consensus reconstruction particularly challenging.
+This LSU assembly dataset (EMPIAR-10076) contains substantial compositional and conformational heterogeneity, making consensus reconstruction particularly challenging.
 Both FSC analysis and visualization show that cryoSplat is more resilient under such conditions, achieving a resolution of $3.30$ Å with fewer artifacts than voxel-based or neural methods, as indicated by the red circle.
-EMPIAR-10180.
-The spliceosome dataset features large-scale motions of the SF3b indicated by the red circle, making consensus reconstruction particularly challenging.
+The spliceosome dataset (EMPIAR-10180) features large-scale motions of the SF3b indicated by the red circle, making consensus reconstruction particularly challenging.
 The reconstructions from backprojection and cryoDRGN show pronounced artifacts in this region, while cryoSplat is more robust to such motions and achieves a resolution of $4.26$ Å.
 FSC analysis further confirms that cryoSplat significantly outperforms the baselines across the frequency range.
-CryoSplat consistently converges within just $5$ epochs, with FSC curves from the $4$th and $5$th epochs tightly overlapping, indicating stable optimization and improved generalization in consensus reconstruction.
-Ablation Studies
+CryoSplat consistently converges within 5 epochs, with FSC curves from the 4th and 5th epochs tightly overlapping, indicating stable optimization and improved generalization.
+4.3 Ablation studies
+This section reports ablation studies of our approach. More results can be found in Appendix E.
 Number of Gaussians.
@@
 Overall, we find that using 10,00010{,}000 Gaussians is sufficient to reconstruct high-resolution volumes that consistently outperform baseline methods across most datasets.
-Runtime efficiency.
+Associated qualitative comparisons are provided in Appendix E.
 […]
-Conclusion
+5 Conclusion
 We present cryoSplat, a novel GMM-based framework that integrates Gaussian splatting with the physics of cryo-EM image formation.
@@
 Appendix
-Appendix A Dataset Details
-We provide detailed statistics and characteristics of the four cryo-EM datasets used in our experiments:
-  • EMPIAR-10028 (Plasmodium falciparum 80S (Pf80S) ribosome) (Wong et al.
-2014): 105,247105{,}247 particle images of size $360\times 360$ pixels at a sampling rate of $1.34$ Å/pixel.
+Appendix A Details of method
+A.1 Real space reconstruction
+According to the Fourier slice theorem (Bracewell, 1956), illustrated in Fig.
+1(c), the 2D Fourier transform of a projection corresponds to a central slice of the 3D Fourier transform of the volume, orthogonal to the projection direction and passing through the origin.
+Based on this property, an alternative and widely adopted formulation models reconstruction directly in the Fourier domain, where the image formation model becomes:
+$\widehat{Y}(k_{x},k_{y})=\widehat{H}(k_{x},k_{y})\cdot\widehat{V}({\bm{W}}^{\top}{\bm{k}})\cdot e^{-2\pi i{\bm{k}}^{\top}{\bm{t}}}+\widehat{{\textnormal{$\epsilon$}}}$, (10)
+where ${\bm{k}}=[k_{x},k_{y},0]^{\top}$ denotes the Cartesian coordinates in Fourier space, and the 2D spectrum $\widehat{Y}$, the CTF $\widehat{H}$ and the 3D spectrum $\widehat{V}$ denote the Fourier transform of $Y$, $H$ and $V$, respectively.
+The noise term $\widehat{{\textnormal{$\epsilon$}}}$ is similarly modeled as independent, zero-mean Gaussian noise in the Fourier domain.
+In this work, departing from most existing approaches that adopt Eq. (10), we instead build our pipeline on Eq. (1), performing homogeneous reconstruction directly in real space.
+A key reason we choose to operate in real space is that it allows us to fully exploit the fast rasterization strategy from 3DGS.
+In high-resolution reconstructions, individual Gaussians in real space have small spatial scales and affect only a few nearby tiles, as shown in Fig.
+5(a).
+This locality means that each GPU thread is responsible for a single pixel and only needs to process a small subset of all Gaussians.
+In contrast, Gaussians in Fourier space become broad as resolution increases, leading to near-global support.
+As a result, each pixel in the frequency domain must aggregate contributions from nearly all Gaussians, making fast rendering impractical.
+Figure 5: Rasterization details.
+(a) A Gaussian with small spatial scales in real space during high-resolution reconstruction overlaps at most four tiles, while in Fourier space it exhibits nearly global support.
+Tile boundaries are indicated by lines.
+(b) For a $4\times 4$ image, the origin of the continuous coordinate system during rasterization is defined differently: FFT-based coordinates place it at $2,2]^{\top}$, whereas image-centered coordinates place it at $1.5,1.5]^{\top}$.
+The origin is marked by a dot.
+Image-centered coordinates induce a phase error of $\pi{\bm{k}}/D$, up to $\pm\pi/2$ at the Nyquist frequency, degrading reconstruction at high frequency.
+The effect diminishes with larger $D$ but never vanishes.
+A.2 Computational details
+Section 3.3.3 describes how a 3D Gaussian is projected along the $z$-axis to form a splat.
+As discussed in Zwicker et al.
+(2002), the splat can be computed analytically by removing the $z$-axis components from the mean and covariance:
+$\begin{cases}\dot{{\bm{\mu}}}=[\dot{\mu}_{x},\dot{\mu}_{y},\dot{\mu}_{z}]^{\top}\Rightarrow[\dot{\mu}_{x},\dot{\mu}_{y}]^{\top}=\tilde{{\bm{\mu}}}\\ \dot{{\bm{\Sigma}}}=\begin{bmatrix}\dot{\sigma}_{xx}&\dot{\sigma}_{xy}&\dot{\sigma}_{xz}\\ \dot{\sigma}_{xy}&\dot{\sigma}_{yy}&\dot{\sigma}_{yz}\\ \dot{\sigma}_{xz}&\dot{\sigma}_{yz}&\dot{\sigma}_{zz}\\ \end{bmatrix}\Rightarrow\begin{bmatrix}\dot{\sigma}_{xx}&\dot{\sigma}_{xy}\\ \dot{\sigma}_{xy}&\dot{\sigma}_{yy}\\ \end{bmatrix}=\tilde{{\bm{\Sigma}}}\end{cases}$ (11)
+thereby enabling an efficient computation of Eq. (7).
+As discussed in Sec.
+3.3.4 and shown in Fig.
+5(b), when a continuous image $X:{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ is rasterized onto pixels ${\bm{X}}\!\in\!{\mathbb{R}}^{D\times D}$, the origin of the continuous coordinate system should be aligned with $\lfloor D/2\rfloor,\lfloor D/2\rfloor]^{\top}$ to match the FFT-based coordinate convention used in cryo-EM.
+Formally,
+${X}_{i,j}=X\left((j-\lfloor\frac{D}{2}\rfloor)\frac{2E}{D},-(i-\lfloor\frac{D}{2}\rfloor)\frac{2E}{D}\right),$ (12)
+where ${X}_{i,j}$ denotes $(i,j)$-th entry of matrix $\bm{X}$}.
+Note that the row and column indices correspond to the $y$- and $x$-axes, respectively, with the $y$-axis flipped during this rasterization.
+In practice, since the PSF corresponds to a large convolution kernel, we apply the contrast transfer function (CTF) in the Fourier domain after rasterization for efficiency:
+${X}_{i,j}=\mathcal{F}^{-1}\left(\widehat{{\bm{H}}}\odot\mathcal{F}\left(\sum_{i=1}^{N}A_{i}\tilde{G}_{i}\left((j-\lfloor\tfrac{D}{2}\rfloor)\tfrac{2E}{D},-(i-\lfloor\tfrac{D}{2}\rfloor)\tfrac{2E}{D}\right)\right)\right),$ (13)
+where $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the Fourier and inverse Fourier transform operators, respectively.
+$\widehat{{\bm{H}}}$ is the rasterized CTF $\widehat{H}$ and $\odot$ denotes element-wise (Hadamard) product.
+Before deriving the gradients, we first define
+[…]
+which is the pre-rasterization continuous image.
+For clarity, we omit the Gaussian index $i$ in the following derivations, as the gradients are computed independently for each Gaussian.
+We denote by $\mathbb{P}$} the set of 2D coordinates corresponding to the centers of rasterized pixels.
+When $\tilde{{\bm{r}}}\in{\mathbb{P}}$, the coordinate $\tilde{{\bm{r}}}=[r_{x},r_{y}]^{\top}$ refers to a discrete sampling location in the image plane.
+The gradients used in the backward pass can be summarized as
+$\left\{\begin{aligned} \frac{\partial\mathcal{L}}{\partial A}&=\sum_{\tilde{{\bm{r}}}\in{\mathbb{P}}}\frac{\partial\mathcal{L}}{\partial Q(\tilde{{\bm{r}}})}\frac{\partial Q(\tilde{{\bm{r}}})}{\partial A}\\ \nabla_{\bm{\mu}}\mathcal{L}&=\sum_{\tilde{{\bm{r}}}\in{\mathbb{P}}}\frac{\partial\mathcal{L}}{\partial Q(\tilde{{\bm{r}}})}\nabla_{\bm{\mu}}Q(\tilde{{\bm{r}}})\\ \nabla_{\bm{s}}\mathcal{L}&=\sum_{\tilde{{\bm{r}}}\in{\mathbb{P}}}\frac{\partial\mathcal{L}}{\partial Q(\tilde{{\bm{r}}})}\nabla_{\bm{\Sigma}}Q(\tilde{{\bm{r}}})\circ\frac{\partial{\bm{\Sigma}}}{\partial{\bm{s}}}\\ \nabla_{\bm{q}}\mathcal{L}&=\sum_{\tilde{{\bm{r}}}\in{\mathbb{P}}}\frac{\partial\mathcal{L}}{\partial Q(\tilde{{\bm{r}}})}\nabla_{\bm{\Sigma}}Q(\tilde{{\bm{r}}})\circ\frac{\partial{\bm{\Sigma}}}{\partial{\bm{q}}}\end{aligned}\right$.
+(15)
+where $\circ$ denotes the composition of Jacobian operators (chain rule). The derivation of gradients with respect to the amplitude $A$ and mean $\bm{\mu}$} is trivial, which can be given directly by
+[…]
+and
+$\left\{\begin{aligned} \nabla_{\tilde{{\bm{\mu}}}}Q&=A\tilde{G}(\tilde{{\bm{r}}})\tilde{{\bm{\Sigma}}}^{-1}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})\\ \nabla_{\bm{\mu}}Q&=W[\nabla_{\tilde{{\bm{\mu}}}}Q^{\top}\;0]^{\top}\end{aligned}\right$.
+(17)
+[…]
+For completeness, we provide the derivation of the covariance gradients $\nabla_{\bm{\Sigma}}Q$, noting that our formulation retains the normalization term, which is omitted in 3DGS (Kerbl et al., 2023).
+Remember
+[…]
+$).\displaystyle=\frac{|\tilde{{\bm{\Sigma}}}^{-1}|^{\frac{1}{2}}}{2\pi}\exp(-\frac{1}{2}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})^{\top}\tilde{{\bm{\Sigma}}}^{-1}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}$)).
+We can first compute
+$\nabla_{\tilde{{\bm{\Sigma}}}^{-1}}Q$ $=A\exp(-\frac{1}{2}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})^{\top}\tilde{{\bm{\Sigma}}}^{-1}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}))\frac{1}{4\pi}|\tilde{{\bm{\Sigma}}}^{-1}|^{-\frac{1}{2}}|\tilde{{\bm{\Sigma}}}^{-1}|\tilde{{\bm{\Sigma}}}^{\top}$ (19)
+$)\displaystyle+A\frac{|\tilde{{\bm{\Sigma}}}^{-1}|^{\frac{1}{2}}}{2\pi}\exp(-\frac{1}{2}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})^{\top}\tilde{{\bm{\Sigma}}}^{-1}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}))(-\tfrac{1}{2}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}$$^{\top})$
+$),\displaystyle=\tfrac{1}{2}A\tilde{G}(\tilde{{\bm{r}}})(\tilde{{\bm{\Sigma}}}-(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}$$^{\top})$,
+and then $\nabla_{\tilde{{\bm{\Sigma}}}}Q=-\tilde{{\bm{\Sigma}}}^{-\top}\nabla_{\tilde{{\bm{\Sigma}}}^{-1}}Q\tilde{{\bm{\Sigma}}}^{-\top}$. Finally,
+$\nabla_{\dot{{\bm{\Sigma}}}}Q=\begin{bmatrix}\nabla_{\tilde{{\bm{\Sigma}}}}Q&{\bm{0}}\\[6.0pt] {\bm{0}}^{\top}&0\end{bmatrix}$. (20)
+The subsequent derivations of $\nabla_{\bm{s}}\mathcal{L}$ and $\nabla_{\bm{q}}\mathcal{L}$ follow exactly the formulation in Kerbl et al. (2023).
+Appendix B Dataset details
+We provide detailed statistics and characteristics of the cryo-EM datasets used in our experiments:
+  • EMPIAR-10028 (Plasmodium falciparum 80S (Pf80S) ribosome) (Wong et al., 2014): 105,247105{,}247 particle images of size $360\times 360$ pixels at a sampling rate of $1.34$ Å/pixel.
 This is a widely used benchmark with high-contrast images and a static structure.
-  • EMPIAR-10049 (RAG1-RAG2 complex) (Ru et al.
-2015): 108,544108{,}544 particles of size $192\times 192$ pixels at $1.23$ Å/pixel.
+  • EMPIAR-10049 (RAG1-RAG2 complex) (Ru et al., 2015): 108,544108{,}544 particles of size $192\times 192$ pixels at $1.23$ Å/pixel.
 This dataset is considered more challenging due to its lower contrast and flexibility in some regions.
   • EMPIAR-10076 (E.
-coli large ribosomal subunit undergoing (LSU) assembly) (Davis et al.
-2016): 131,899131{,}899 particles of size $320\times 320$ pixels at $1.31$ Å/pixel.
+coli large ribosomal subunit undergoing (LSU) assembly) (Davis et al., 2016): 131,899131{,}899 particles of size $320\times 320$ pixels at $1.31$ Å/pixel.
 This dataset contains substantial conformational and compositional heterogeneity, which poses a challenge to homogeneous modeling.
-  • EMPIAR-10180 (Pre-catalytic spliceosome) (Plaschka, Lin, and Nagai 2017): 327,490327{,}490 particles of size $320\times 320$ pixels at $1.69$ Å/pixel.
+  • EMPIAR-10180 (Pre-catalytic spliceosome) (Plaschka et al., 2017): 327,490327{,}490 particles of size $320\times 320$ pixels at $1.69$ Å/pixel.
 It samples a continuum of conformations, particularly involving large-scale motions of the SF3b subcomplex.
-Appendix B Fourier Shell Correlation
-To evaluate reconstruction quality on real datasets where ground truth volumes are unavailable, we adopt the gold standard Fourier Shell Correlation (FSC) (Van Heel and Schatz 2005), following established protocols.
+  • Synthetic 80S ribosome: We construct a synthetic dataset of the 80S ribosome with 100,000100{,}000 particles using Relion (Scheres, 2016a), following the protocol of Levy et al.
+(2022a).
+The electron scattering potential is derived in ChimeraX (Pettersen et al., 2021) at a resolution of $6.0$ Å/pixel, based on two atomic models: the small subunit (PDB 3J7A) and the large subunit (PDB 3J79) (Wong et al., 2014).
+Each particle image is $128\times 128$ pixels with a pixel size of $3.77$ Å/pixel.
+Orientations are uniformly sampled over $\mathrm{SO}(3)$, and all images are centered without translations.
+Defocus values for the CTF are randomly drawn from log-normal distributions following Levy et al.
+(2022a), and zero-mean white Gaussian noise with varying signal-to-noise ratios (SNRs) is added.
+Appendix C Fourier shell correlation
+To evaluate reconstruction quality on real datasets without ground truth volumes, we adopt the gold standard Fourier Shell Correlation (FSC) (Van Heel & Schatz, 2005), following established protocols.
 Each dataset is randomly split into two halves, and the reconstruction algorithm is applied independently to each subset.
-Let the resulting volumes be $\widehat{\mathcal{V}}_{a}(\mathbf{k})$ and $\widehat{\mathcal{V}}_{b}(\mathbf{k})$, representing their Fourier transforms.
-The FSC is computed as a function of spatial frequency $k=|\mathbf{k}|$ using the following formula:
-$\mathrm{FSC}(k)=\frac{\sum\limits_{|\mathbf{k}|=k}\widehat{\mathcal{V}}_{a}(\mathbf{k})\cdot\widehat{\mathcal{V}}_{b}(\mathbf{k})^{*}}{\sqrt{\Big{(}\sum\limits_{|\mathbf{k}|=k}|\widehat{\mathcal{V}}_{a}(\mathbf{k})|^{2}\Big{)}\Big{(}\sum\limits_{|\mathbf{k}|=k}|\widehat{\mathcal{V}}_{b}(\mathbf{k})|^{2}\Big{)}}}.$ (B.1)
+Let the resulting volumes be $\widehat{V}_{a}({\bm{k}})$ and $\widehat{V}_{b}({\bm{k}})$, representing their Fourier transforms.
+The FSC is computed as a function of frequency $k$ using the following formula:
+$\mathrm{FSC}(k)=\frac{\sum\limits_{\|{\bm{k}}\|_{2}=k}\widehat{V}_{a}({\bm{k}})\cdot\widehat{V}_{b}({\bm{k}})^{*}}{\sqrt{\Big(\sum\limits_{\|{\bm{k}}\|_{2}=k}|\widehat{V}_{a}({\bm{k}})|^{2}\Big)\Big(\sum\limits_{\|{\bm{k}}\|_{2}=k}|\widehat{V}_{b}({\bm{k}})|^{2}\Big)}}.$ (21)
 This metric quantifies the correlation between two independently reconstructed volumes within concentric shells in Fourier space.
 The spatial resolution is defined as the frequency where the FSC curve drops below the $0.143$ threshold, indicating the limit of reproducible structural detail.
-Appendix C More Implementation Details
-For all experiments, we do not apply any windowing to the observed particle images during preprocessing.
-The 3D volume is defined over the domain $E,E]^{3}$, and each 2D projection is assumed to lie within $E,E]^{2}$ in the image plane, where $E=0.5$ defines the spatial extent.
-Intuition Behind Initialization
-These values are fixed but grounded in straightforward statistical intuition.
-We observe that most particles are concentrated within a spherical region of radius $E/2$.
+Appendix D More implementation details
+D.1 Intuition behind initialization
+The values used in initialization are fixed but grounded in straightforward statistical intuition.
+We observe that most particles are concentrated within a spherical region of radius $E/2$, where $E=0.5$ defines the spatial extent as mentioned in Sec.
+4.1.
 To reflect this prior and accelerate convergence, we initialize the Gaussian means within this region.
@@
 Finally, to maintain consistent overall energy across varying numbers of Gaussians, we initialize the amplitude as $A=1/(2N)$, where $N$ is the total number of Gaussians.
+D.2 Intuition behind learning rate
+In the original 3DGS (Kerbl et al., 2023), different learning rates are assigned to different types of Gaussian parameters (means, scales, rotations, opacities).
+While this works well in novel view synthesis, it introduces instability in cryo-EM reconstruction.
+Let the full parameter vector be $\bm{\theta}=[\mu_{x},\mu_{y},\mu_{z},s_{x},s_{y},s_{z},q_{w},q_{x},q_{y},q_{z},A]^{\top}$.
+In gradient descent optimization, the direction of parameter updates is determined by the gradient $\nabla_{\bm{\theta}}\mathcal{L}$.
+Unequal learning rates distort this direction by scaling different components unequally, which can lead to divergence.
+We observe that such practice causes Gaussians to spread uncontrollably in early iterations and finally diverge.
+To avoid this, we adopt a single unified learning rate across all parameter types, preserving the intended descent direction and ensuring stable convergence.
+D.3 Optimization algorithm
+Our optimization algorithm is summarized in Algorithm 1.
+Unlike Kerbl et al.
+(2023), which uses gradient magnitude as the criterion for splitting and cloning Gaussians, we observe that gradients are not a reliable indicator for densification in cryo-EM reconstruction.
+Furthermore, elaborate densification schemes are generally unnecessary, as our method seldom suffers from significant local minima owing to its close consistency with cryo-EM imaging physics.
+Nevertheless, we retain a simple densification option to balance efficiency and resolution: fewer Gaussians enable faster training, whereas more Gaussians yield higher-resolution reconstructions, as demonstrated in Fig.
+4.
+Algorithm 1 Optimization and Densification
+NN: number of Gaussians
+$D$: side length of the observed particle images
+   $\bm{\Theta}$$\leftarrow$ InitAttributes(NN) $\triangleright$ Positions, Scales, Quaternions, Amplitudes
+   $i\leftarrow$ 0 $\triangleright$ Epoch Count
+   while not converged do
+    for ($\bm{Y}$}, $\bm{W}$}, $\bm{t}$}, $\widehat{{\bm{H}}}$) in Dataloader() do $\triangleright$ Observed Image, Rotation, Translation, CTF
+     $\bm{Y}$$\leftarrow$ FourierShift($\bm{Y}$}, $\bm{t}$}) $\triangleright$ Center Alignment
+     $\bm{Q}$$\leftarrow$ Rasterize($\bm{\Theta}$}, $\bm{W}$}, $D$) $\triangleright$ Algorithm 2
+     $\bm{X}$$\leftarrow$ ApplyCTF($\bm{Q}$}, $\widehat{{\bm{H}}}$) $\triangleright$ Apply CTF
+     $\mathcal{L}\leftarrow$ Loss($\bm{X}$}, $\bm{Y}$}) $\triangleright$ Loss
+     $\bm{\Theta}$$\leftarrow$ Adam($\nabla\mathcal{L}$) $\triangleright$ Backprop and Step
+    end for
+    if IsDoubleGaussians($i$) then $\triangleright$ (Optional) Densification
+     for all Gaussian($\bm{\mu}$}, $\bm{s}$}, $\bm{q}$}, AA) in $\bm{\Theta}$} do
+      SplitGaussian($\bm{\mu}$}, $\bm{s}$}, $\bm{q}$}, AA)
+     end for
+    end if
+    $i\leftarrow i+1$
+   end while
+Algorithm 2 CUDA-accelerated Rasterization
+$\bm{\Theta}$}: Gaussian parameters
+$\bm{W}$}: viewing transformation matrix
+$D$: side length of the observed particle images
+   function Rasterize($\bm{\Theta}$}, $\bm{W}$}, $D$)
+    $\bm{\mu}$}, $\bm{\Sigma}$}, $A\leftarrow$ BuildGaussians($\bm{\Theta}$})
+    $\dot{{\bm{\mu}}}$, $\dot{{\bm{\Sigma}}}\leftarrow$ ViewingTransform($\bm{\mu}$}, $\bm{\Sigma}$}, $\bm{W}$}) $\triangleright$ Viewing Transformation
+    $\tilde{{\bm{\mu}}}$, $\tilde{{\bm{\Sigma}}}\leftarrow$ Projection($\dot{{\bm{\mu}}}$, $\dot{{\bm{\Sigma}}}$) $\triangleright$ Orthogonal Projection
+    $T\leftarrow$ CreateTiles($D$) $\triangleright$ Tile Count
+    $\bm{L}$}, $\bm{K}$$\leftarrow$ DuplicateWithKeys($\tilde{{\bm{\mu}}}$, $T$) $\triangleright$ List of Indices and Keys
+    SortByKeys($\bm{K}$}, $\bm{L}$})
+    $\bm{R}$$\leftarrow$ IdentifyTileRanges($T$, $\bm{K}$})
+    ${\bm{Q}}\leftarrow{\bm{0}}$ $\triangleright$ Init Canvas
+    for all Tile $\bm{t}$} in $\bm{Q}$} do
+     for all Pixel $\bm{p}$} in $\bm{t}$} do
+      $\bm{r}$$\leftarrow$ GetTileRange($\bm{R}$}, $\bm{t}$})
+      ${\bm{Q}}({\bm{p}})\leftarrow$ SumSplats($\bm{p}$}, $\bm{L}$}, $\bm{r}$}, $\bm{K}$}, $\tilde{{\bm{\mu}}}$, $\tilde{{\bm{\Sigma}}}$, $A$)
+     end for
+    end for
+    return $\bm{Q}$}
+   end function
+D.4 Details of the rasterizer
+The details of the rasterizer are summarized in Algorithm 2.
+We follow the tile-based rasterization framework of Kerbl et al.
+(2023), where the output image is divided into $16\times 16$ pixel tiles, and each splat is instantiated in every tile it overlaps.
+The splat instances are then assigned keys for sorting, after which each tile can be processed efficiently by locating the corresponding ranges in the sorted list.
+Since pixels are computed in parallel, the runtime is primarily determined by the maximum number of Gaussians within any tile.
+For more details, we refer the reader to Kerbl et al.
+(2023).
+Appendix E Additional results
+E.1 Multimedia results
+We provide videos of rotating reconstructed volumes in the supplementary material for all methods and datasets to facilitate visual comparison.
+Table 1: GPU memory usage across reconstruction methods at resolutions ($D=192$, $D=256$).
 Methods # Params Settings Batch Size GPU Mem. (MiB)
 $D=192$ $D=256$
-Backprojection $(D+1)^{3}$ 11 $508$ $396$
-CryoDRGN (2021a) $(6\cdot\lfloor D/2\rfloor+L+3)\cdot C$ C=1,024C=1{,}024 11 $680$ 1,0081{,}008
-$L\cdot C^{2}+2$ $L=3$ 88 2,5602{,}560 4,9064{,}906
-CryoSplat (Ours) $11\cdot N$ N=2,048N=2{,}048 11 $344$ $346$
+Backprojection $(D+1)^{3}$ 1 508 396
+CryoDRGN (Zhong et al., 2021a) $(6\cdot\lfloor D/2\rfloor+L+3)\cdot C$ C=1,024C=1{,}024 1 680 1,008
+$L\cdot C^{2}+2$ $L=3$ 8 2,560 4,906
+CryoSplat (Ours) $11\cdot N$ N=2,048N=2{,}048 1 344 346
 N=3,072N=3{,}072 344 348
@@
 N=30, [^0] N=30{,}000 376 378
-Table S1: Comparison of GPU memory usage across reconstruction methods at $D=192$ and $D=256$.
-CryoSplat achieves the lowest and most stable memory footprint, while cryoDRGN incurs significantly higher usage due to its deep neural architecture.
-All measurements were performed on a single NVIDIA RTX 3090, reflecting the memory usage during the reconstruction of a single full density map.
-Memory for voxelization and FSC-related half-map reconstruction is not included.
-Intuition Behind Learning Rate
-In the original 3DGS (Kerbl et al.
-2023), different learning rates are assigned to different types of Gaussian parameters (means, scales, rotations, opacities).
-While this works well in novel view synthesis, it introduces instability in cryo-EM reconstruction.
-Let the full parameter vector be $\boldsymbol{\theta}=[\mu_{x},\mu_{y},\mu_{z},s_{x},s_{y},s_{z},q_{w},q_{x},q_{y},q_{z},A]^{\top}$.
-In gradient descent optimization, the direction of parameter updates is determined by the gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}$.
-Unequal learning rates distort this direction by scaling different components unequally, which can lead to divergence.
-We observe that such practice causes Gaussians to spread uncontrollably in early iterations and finally diverge.
-To avoid this, we adopt a single unified learning rate across all parameter types, preserving the intended descent direction and ensuring stable convergence.
-Appendix D Additional Results
-Memory Usage
+E.2 Memory usage
 Tab.
-S1 presents a comparison of GPU memory usage across different reconstruction methods.
-CryoDRGN (Zhong et al.
-2021a) exhibits the highest memory footprint, exceeding $2.5$ GiB at $D=192$ and approaching $5$ GiB at $D=256$, primarily due to its deep neural decoder and a larger batch size of 8.
+1 compares GPU memory usage across different reconstruction methods.
+CryoDRGN (Zhong et al., 2021a) exhibits the highest memory footprint, exceeding $2.5$ GiB at $D=192$ and approaching $5$ GiB at $D=256$, primarily due to its deep neural decoder and a larger batch size of 8.
 Interestingly, backprojection consumes more memory at $D=192$ than at $D=256$, which may be attributed to implementation-specific factors such as padding overhead or kernel-level optimizations that favor power-of-two dimensions.
@@
 In contrast, CryoSplat demonstrates consistently low memory usage across all configurations.
-Even with as many as 30,00030{,}000 Gaussians, cryoSplat maintains a GPU memory footprint below $380$ MiB, with negligible variation across resolutions.
+Even with as many as 30,00030{,}000 Gaussians, CryoSplat maintains a memory footprint below $380$ MiB, with negligible variation across resolutions.
 This efficiency underscores the scalability and suitability of CryoSplat for large-scale or memory-constrained cryo-EM reconstruction scenarios.
-Figure S1: Qualitative and quantitative comparison of isotropic and anisotropic GMMs (N=30, [^0] N=30{,}000) on four real datasets.
-FSC curves show that anisotropic Gaussians consistently achieve higher correlations across spatial frequencies, indicating improved reconstruction accuracy and robustness.
-Volume visualizations further reveal that anisotropic GMMs better recover fine structural details and directionally extended shapes, whereas isotropic Gaussians tend to fragment such regions.
-Isotropic vs. Anisotropic
-We investigate how the choice between isotropic and anisotropic Gaussians affects reconstruction quality.
-While cryoSplat adopts anisotropic Gaussians to represent 3D volumes, it remains compatible with the isotropic formulation commonly used in prior work (2021; 2023; 2023; 2024; 2025).
+Figure 6: Qualitative and quantitative comparison of isotropic and anisotropic GMMs (N=30, [^0] N=30{,}000) on two datasets (EMPIAR-10028 and EMPIAR-10049).
+FSC curves show that anisotropic Gaussians consistently achieve higher correlations across spatial frequencies, indicating improved reconstruction accuracy.
+Volume visualizations further reveal that anisotropic GMMs better recover fine structural details and elongated features, whereas isotropic Gaussians tend to fragment such regions.
+E.3 Isotropic vs. anisotropic
+CryoSplat represents 3D volumes using anisotropic Gaussians while remaining fully compatible with the isotropic formulation widely adopted in prior works (Chen & Ludtke, 2021; Chen et al., 2023a; b; Schwab et al., 2024; Chen, 2025).
+When the scaling is isotropic, $s_{x}=s_{y}=s_{z}=\sigma$, the anisotropic Gaussian exactly reduces to the standard isotropic form:
+$G({\bm{r}}|{\bm{\mu}},\sigma)=\frac{1}{(2\pi)^{\frac{3}{2}}\sigma^{3}}\exp(-\frac{\|{\bm{r}}-{\bm{\mu}}\|_{2}^{2}}{2\sigma^{2}}),$ (22)
+allowing direct integration into existing isotropic GMM-based pipelines.
+We investigate the impact of isotropic versus anisotropic Gaussians on reconstruction quality.
 As shown in Fig.
-S1, anisotropic GMMs achieve higher FSC scores across spatial frequencies and produce sharper, more detailed structures.
-Subjectively, isotropic Gaussians struggle to capture directionally extended structures and are more prone to noise, often resulting in oversmoothing or distortion.
-This geometric rigidity may impair their ability to fit complex structures during early optimization, contributing to the unstable convergence observed with random initialization in prior methods.
+6, anisotropic GMMs achieve higher FSC scores across spatial frequencies and produce sharper, more detailed structures.
+Subjectively, isotropic Gaussians struggle to capture elongated features and are often captured by noise, which may contribute to the unstable convergence from random initialization reported in previous methods.
 These results highlight the improved representational capacity and reconstruction robustness enabled by anisotropic modeling.
-Figure S2: Qualitative evaluation of reconstruction performance with different numbers of Gaussians.
+E.4 Number of Gaussians
+Figure 7: Qualitative evaluation of reconstruction performance with different numbers of Gaussians.
 Increasing the number of Gaussians leads to visibly improved reconstructions, with finer structural details and enhanced sharpness.
 Red arrows mark representative regions that highlight the qualitative differences for clearer comparison across settings.
-Number of Gaussians
 We present visual comparisons of reconstruction results using different numbers of Gaussians.
 As shown in Fig.
-S2, increasing the number of components yields progressively sharper and more detailed structures.
-These qualitative observations align with the quantitative improvements in FSC curves reported in the main paper.
+7, increasing the number of components yields progressively sharper and more detailed structures.
+These qualitative observations align with the quantitative improvements in FSC curves reported in Fig.
+3.
 Red arrows highlight representative regions where the differences in reconstruction quality are especially pronounced, facilitating direct visual comparison across settings.
+E.5 Signal-to-noise ratio
+Figure 8: Reconstruction performance under varying SNRs.
+(top) Ground-truth (GT) and reconstructed volumes at different SNR levels.
+(bottom) Example synthetic particle images corresponding to each SNR.
+(right) FSC curves between GT and reconstructed volumes across SNRs.
+We study the effect of SNR levels on cryoSplat with 5,1205{,}120 Gaussians using the synthetic 80S dataset described in Sec.
+B.
+Figure 8 shows example synthetic particles, reconstructed volumes, and FSC curves under varying SNRs.
+FSCs are computed between the ground truth (GT) and reconstructed volumes.
+Overall, cryoSplat shows strong noise robustness: SNRs above 0 dB have little impact on reconstruction; high resolution is preserved even under severe noise at 15-15 dB, and reconstructions remain satisfactory at 20-20 dB, despite particles being barely visible.